### v12

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
The Seymour Second Neighborhood Conjecture (SSNC) claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we establish the Graph Level Order (GLOVER) data structure, which orders the nodes by shortest path from a minimum out-degree node and establishes a well-ordering of rooted neighborhoods. This data structure allows for the construction of decreasing sequences of subsets of nodes and allows us to partition transitive triangles into distinct sets. The decreasing sequence of nodes shows the non-existence of counterexamples to the SSNC and precisely identifies a path to the required node. Further, our algorithmic approach finds the occurrence of dense graphs inside the rooted neighborhoods. Beyond theoretical implications, the algorithm and data structure have practical applications in data science, network optimization and algorithm design.
Keywords: data structure, decreasing neighborhood sequence property, rooted neighborhood, exterior neighbor, minimum degree node, transitive triangle, load balancing
1 Introduction
The Seymour Second Neighborhood Conjecture (SSNC), proposed by Paul Seymour in 1990, is a deceptively easy to state research problem, but one that has shown much resistance to proof. It asks whether every oriented graph contains a vertex whose second out-neighborhood is at least as large as its first. We provide an affirmative answer to this long-standing open problem. Our resolution hinges on the development of a powerful data structure, the Graph Level Order (GLOVER), which organizes the vertices of an oriented graph into a hierarchy that produces an inherent doubling behavior in neighborhood sizes. This data structure not only allows for a constructive proof of the conjecture but also yields an efficient algorithm for explicitly identifying such a vertex.
Unlike previous approaches, this framework allows us to locate vertices that satisfy the SSNC explicitly by way of a path to those vertices. The development of this data structure has been an interdisciplinary challenge, which drew insights from combinatorics, algorithm design, and network theory. With a background in computer programming and data science, our search was for not only provable theorems, but usable and verifiable algorithms. Even with the abstract concepts we produced via group theory, we have tried to maintain algorithmic accessibility throughout.
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G$, there exists a vertex $v\in G$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$, respectively.
This conjecture was first published by Nathaniel Dean and Brenda Latka in 1995 [8], where they proposed a related version specific to tournaments, which are complete oriented graphs. The tournament case was proved by Fisher [11] in 1996, with an alternative proof later provided by Havet and Thomassé [18] in 2000.
The work on the solutions to the Dean Tournament Conjecture is very important in the study of the second out-neighborhood conjecture. Fisher’s approach [11] did not yield a method for locating a specific vertex, but it did prove that such a vertex must exist. The method did so with probability, by showing that there exists at least one node whose second out-neighborhood must be at least as large as its first. Later, Havet and Thomassé [18] advanced the problem further by using median orders to provide a more constructive approach. This technique identifies a candidate vertex and a path to that vertex. Median orders is an NP-hard problem and introduces computational complexities to their approach. Still, this was a significant breakthrough in combinatorics, particularly for tournament theory, as median-orders were now available for usage.
Kaneko and Locke [20] showed that the SSNC holds for graphs with a minimum degree of at most six. This work provided a ground level for a lot of the research on the SSNC problem. However, it also brought another NP-Hard problem, dominating sets, into the realm of the SSNC. Unfortunately, for cases greater than six, the complexity was too great to solve efficiently.
Chen et al. [5] established a lower bound for the SSNC, showing that some vertex satisfies a relaxed version of the condition. This work was extended by Huang and F. Peng [17], who incorporated third neighborhoods and improved bounds on the fraction of Seymour vertices using roots of polynomial inequalities. Later they were able to improve the bounds on the number of Seymour vertices in a graph up to a fraction $\gamma$, where $\gamma$ is the real root in the range [0,1] of the equation $08x^{5}+4x^{4}-12x^{3}-7x^{2}+2x+4=$ 0.
These results originate from Constraint Satisfaction Programming (CSP). This is a method that tries to frame the SSNC as a series of logical constraints to be satisfied and optimized by way of SAT solvers. This direction has led to increasingly tighter bounds, but it again operates within the realm of the NP-hard problem CSP. This makes both scalability and generalization a challenge. Also, resolving SSNC through CSP offers little insight into the graphs themselves, which is an issue we aim to resolve with our constructive algorithmic approach.
In recent work, Diaz et al. [9] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [1], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC. The study by Diaz et al. examines randomly generated graphs, specifically using binomial random graphs with random orientations. It demonstrates that in these constructions there are typically Seymour vertices. These findings are valuable for understanding the average-case behavior of oriented graphs, and this is a major step towards resolving the conjecture. However, identifying Seymour vertices in randomly generated cases fails to identify properties of graphs that either guarantee or prevent the conjecture’s truth. There are many fundamental set-theoretic and graph-theoretic questions that remain open. These are questions that concern the connectivity of graphs, degree distribution, and neighborhood growth.
Nevertheless, the Diaz et al. [9] paper does make meaningful progress by considering the contrapositive of the Seymour conjecture. In particular, Propositions 4 and 5 in the paper explore the implications of the conjecture being false. They show that in such a case, there would be infinitely many strongly connected oriented graphs without Seymour vertices, even if we had a bounded minimum out-degree. This line of reasoning draws resemblance to the Decreasing Neighborhood Sequence Property (DNSP) introduced in this paper. Though not mentioned in their work and having different implications, the counterexample modeling echoes the motivations that led to the development of DNSP here.
Building upon this understanding of local structures, our work examines a particularly insightful contribution that comes from Brantner et al. [2], who examined the role of specific subgraph structures in relation to the SSNC. Their work introduced and formalized two key patterns: transitive triangles and Seymour diamonds in the context of SSNC. A transitive triangle consists of three nodes and three directed edges, forming a configuration where two edges share a common head—an arrangement that creates a shortcut in the graph’s reachability. A Seymour diamond involves four directed edges connecting four nodes, with two edges originating from a common source and converging again at a shared target. Brantner et al. showed that if a directed graph is free of transitive triangles, then it necessarily contains a Seymour vertex—a node whose second out-neighborhood is at least as large as its first. This result not only builds intuition about the conjecture but also suggests that the presence of certain local configurations might be obstructive to the existence of a Seymour vertex. Our work examines how larger, layered neighborhood interactions—particularly in the presence of overlapping or nested transitive triangle substructures—can be systematically analyzed through a new data structure.
The concepts of $m$-free, $k$-transitive, and $k$-anti transitive graphs are also relevant to this paper [6]. A directed graph $D$ is called $m$-free if it contains no directed cycles of length at most $m$. Daamouch also considered $k$-transitive graphs, which are graphs that had these transitive shortcuts for $k\leq 6$, and $k$-anti transitive graphs, for $k\leq 4$. These are graphs without these shortcuts. Daamouch related these extensions of transitivity and anti-transitivity to graphs necessarily having a Seymour vertex. Later, [16] Hassan et al. extended these results to 6-anti transitive graphs.
Transitive triangles are fundamental in the structure of real-world networks, such as those underlying social media and friend recommendation systems. The SSNC is not just a mathematical problem, we are attempting to solve a real world problem. Therefore, any algorithm or framework attempting to resolve the SSNC must be capable of operating meaningfully on graphs that contain them.
The Graph Level Order data structure provides exactly this capability. By organizing nodes based on their distance from a minimum out-degree vertex, it reveals that transitive triangles are not uniform obstacles. Instead, they can be partitioned into distinct cases based on the relative distance of their vertices. This insight is not simply novel, it is necessary. Without formalizing these rooted neighborhoods, critical properties such as decreasing neighborhood size and the influence of back arcs would remain unprovable. The Graph Level Order enables the distinction and handling of cases that were previously inseparable, and is thus essential to the approach developed in this paper.
Variations of the question have asked whether the conjecture holds for graphs lacking certain substructures. For example, Fidler and Yuster [12] proved the conjecture true for tournaments when multiple edges are removed from the same node and orientations of complete graphs missing a star. Later Mniny and Ghazal [21] proved the conjecture true for oriented graphs missing $C_{4},\overline{C_{4}},S_{3}$ chair or co-chair. Following this, Ghazal proved [14] the conjecture true for tournaments missing a star. Daamouch et al. [7] proved the conjecture true for tournaments missing two stars or disjoint paths.
Also of note is the 2021 poster presentation by Illia Nalyvaiko [22], which offered another contra-positive-style approach to explore the SSNC somewhat similar to the Decreasing Neighborhood Sequence Property (DNSP). In his poster, he introduced a penalty function to evaluate the likelihood of a vertex of a counterexample. While this is a different intuition than DNSP, both approaches represent a shift from verifying the conjecture to questioning the possibility of the bounds of the conjecture itself. Given the historical significance of this problem, Nalyvaiko’s contrapositive-style approach offered a novel perspective that resonated with the motivations behind our Decreasing Neighborhood Sequence Property.
Prior investigations of the SSNC have largely focused on entire graphs in search of a Seymour vertex, often neglecting the local conditions under which individual vertices satisfy—or fail to satisfy—the conjecture. Our approach inverts this perspective: we begin with individual vertices and analyze the structural and neighborhood-based features that determine their behavior. This localized view naturally gives rise to new lemmas and insights, ultimately allowing us to constructively identify Seymour vertices and understand the dynamics of their neighborhoods in a principled way.
Rather than simply asking whether a graph satisfies the conjecture by possessing a Seymour vertex, we turn our attention to the nodes that fail to be Seymour vertices. We ask this question not explicitly for a particular node, but for every node in the graph. This enables us to engage in a search, because we are looking for conditions that need to be satisfied for every node in the graph. As such, this proof for a degree doubling node is like building a house of cards. There will be a lot of information that needs to be true for every node to fail to be a degree doubling node. As we continue to try to build this house of cards, what we will see is that it will lead to one of two things: either the house will eventually collapse, leading to a contradiction and producing that degree doubling node, or we finish constructing the house, and every node fails to be a degree doubling node. This later statement would prove the conjecture false.
Beyond its theoretical appeal, the SSNC has practical applications in domains like social networks, epidemiology, algorithm design, and network A/B testing. For example, in disease control, we need to identify individuals whose influence spans two levels of contact. We could target interventions more effectively and mitigate outbreaks. In network-based A/B testing, it is crucial to have disjoint treatment and control groups to avoid interference. That is a fundamental feature of the Graph Level Order data structure. Furthermore, within these communities, the algorithm for SSNC helps us identify highly influential individuals, allowing for more strategic experimental design and more accurate evaluation of interventions.
Unlike the purely existential proof approaches we have considered, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
This paper claims a complete proof of the SSNC using a proof by contradiction-based argument, fundamentally supported by a novel and essential data structure: the Graph Level Order. While the idea of partitioning nodes into neighborhoods is a common technique. Our central findings become evident precisely where BFS ends. We formalize the intricate relationships between neighborhoods, within themselves, to each other, and with the root node. This establishes a powerful framework for understanding the SSNC. The Graph Level Order is not merely a traversal algorithm. It systematically defines the necessary implications for these neighborhoods to uphold the conjecture. For instance, it allows us to decompose transitive triangles, a long-standing barrier to progress on the conjecture, into six distinct, manageable cases. Our approach is constructive and grounded in standard algorithmic tools like partition, cycles, divide and conquer, and traversal and offers linear-time complexity. Beyond merely solving the conjecture, it provides a straightforward blueprint for identifying the node that satisfies it. All key lemmas, theorems and corollaries are all rigorously proven and all terminology strictly adheres to standard graph-theoretic conventions.
This paper begins by presenting the methodology behind our approach. We then define the key terminology and tools required to understand the conjecture. We follow that with lemmas and examples that demonstrate its utility. We then introduce the novel data structure designed to address key challenges, and then we proceed to explore how this data structure can be combined with the Decreasing Neighborhood Sequence Property. We examine the role of back arcs and their role in the conjecture. Our main theorem follows, alongside a detailed analysis of the associated algorithm’s complexity. The paper concludes with potential applications of our approach and insights into future directions for research.
2 Exploring the Contrapositive
Our approach to SSNC was driven by a desire to explore the contrapositive. We were not directly searching for a counterexample. Instead, we were aiming to understand the implications of the conjecture, and what it would mean to assume it false. As data scientists, we naturally gravitated towards an exploratory methodology, programming tools to observe patterns, features, and potential contradictions arising from this assumption. This approach allows us to define concrete conditions that, if violated, directly lead to the identification of a degree-doubling node through our algorithmic process.
One fundamental question we must answer is why we chose proof by contradiction when so few others who have investigated this problem did so. First, from a practical point of view, it is necessary to use all tools at our disposal. Proof by contradiction is one of those tools and is logically equivalent to alternative proof methods, like direct proof. As computer programmers, though, what proof by contradiction does is re-frames the problem from a theoretical point of view into an actionable one. That is, the proof by contradiction gives us an extra assumption that allows us to construct datasets and concrete conditions that can guide the development of an algorithm or insight. Finally, we recognize that fresh perspectives on a long-standing problem are difficult to come by. Investigating the contrapositive offered the insight to explore less-traveled paths and bring about new proofs by simply taking ”The Road Less Traveled”. This offers a higher probability of a breakthrough by simply avoiding the obstacles that previous authors have faced, like complexity issues.
For the Seymour Second Neighborhood Conjecture (SSNC), an additional insight came both from intuition—shaped by our work on the Decreasing Neighborhood Sequence Property (DNSP), which is a direct negation of the conjecture but at the node level—and from a new structural feature: exterior neighbors. While interior neighbors (which are a particular type of transitive triangle) have appeared in the literature in various forms, exterior neighbors have largely gone unnoticed. Yet these two concepts are fundamentally linked. They are, in essence, two sides of the same coin—capturing local and extended connectivity within a graph. The SSNC concerns both first and second out-neighbors, so any attempt to resolve the conjecture rigorously must treat both interior and exterior neighbors in parallel. We cannot fully understand one without the other. Interior neighbors are neighbors shared with a parent from the previous neighborhood within and another member of the same neighborhood. Exterior neighbors, on the other hand, are neighbors between a parent and child in consecutive neighborhoods that are not shared by the parent. This will be defined formally in Section 5 (Graph Level Order). What we will see is that this extra data point allows us to build a traversal algorithm through the graph. These exterior arcs connect to the second neighbors of the parent node, while the interior arcs connect to the first neighbors of the parent node.
2.1 A Programmer’s Insight
Early in the research process, we began to notice that the SSNC places two fundamental yet distinct graph metrics in competition: distance and degree. The conjecture seeks for a relationship between the size of a node’s first out-neighborhood, which is based on direct distance, and its second out-neighborhood, which is based on a distance of two. However, the question the conjecture asks is for a node whose degree doubles. This shows two metrics that are intertwined into one conjecture. Other graph metrics such as centrality or tree width could be considered, but they are not intrinsic to the conjecture. Degree and distance are explicitly stated, and any resolution of the SSNC must contend with their relationship.
When we were faced with the question of which metric to prioritize for our initial exploration and partitioning strategy, both metrics gave reasonable options, but for different reasons. For example, every node has an out-degree, and thus there exists a minimum value of these out-degrees. This is the out-degree value that all other nodes in the graph must at least have. This was a reasonable starting point for exploratory analysis. On the contrary, the distance metric has properties like a total order that could prove useful.
Once we had chosen to start with a minimum out-degree node, our approach remained largely procedural, guided by small, illustrative examples. Before we partitioned the graph, even small graphs offered little insight into the graph’s structure. The main theme of proof by contradiction is that we did not want any node’s degree to double. This is what guided our early exploratory process through the graph. Partitioning the graph by distance from the out-degree of the root node (by $\delta$) was not a goal. Instead, we observed that, much like we will see in Examples 4.1 and 4.2, the nodes were naturally grouping themselves into rooted neighborhoods based on their distance from the anchor node. This partitioning was not imposed, it emerged. This became a foundational insight: by organizing nodes in this manner, into these distance-based layers, we gained a clearer view of their roles and dependencies. The partitioning helped us to understand interior flow, exterior flow and identify patterns. This eventually helped us to formalize a framework in which degree doubling could be analyzed systematically.
The first observation we see from Figure 1 is numerical is that these neighborhoods are decreasing in size. We wondered if that was a coincidence, or a consequence of the graph’s local shape. Or did it hold more broadly. This is a proof by contradiction, so the reason for that in this example is because nodes cannot have their second neighborhoods larger than their first neighborhoods. We called this the Decreasing Neighborhood Sequence Property (DNSP) for this reason. Then we notice something more interesting: the correlation of the interior out-degree (the blue arcs) with the neighborhood distance from the root node. Each node in the first neighborhood has an interior out-degree of one. Then each node in the second neighborhood has an interior out-degree of two. A natural question was arising, are these out-degrees typical, or exceptions? If they were typical, can we leverage them towards finding a degree doubling node?
Another pattern that we noticed was that some nodes had first neighbors within the same neighborhood, sharing a common parent. This formed our concept of interior neighbors. Other nodes had neighbors in the next neighborhood, that were the parent’s second neighbor. These are the nodes we identified as exterior neighbors. This was not just a cosmetic difference. We could partition a node’s degree into interior degree and exterior degree, unlike first neighbors and second neighbors. This helps to uncover the mechanics of load sharing, what is necessary for a node to not be a degree doubling node in a graph. It also was not just present in the first partition, but we were seeing it throughout the graph. Interior neighbors represented the interlocking within the neighborhood, while exterior neighbors represented forward pressure. Recognizing and formalizing this split became essential for understanding how rooted neighborhoods either sustain themselves or collapse under the absence of degree doubling nodes. More importantly were the questions of whether these observations were simple examples, or could they be backed up with mathematical rigor.
Figure 1: This is an early visualization of the SSNC through JavaScript’s Canvas. Nodes are colored based on their distance from the node 00. Interior arcs are blue, while exterior arcs are gray. We see that there are seven nodes in the first rooted neighborhood, matching the degree of node 0. There are six nodes in the second rooted neighborhood, and five nodes in the third rooted neighborhood. Each node in the first rooted neighborhood has degree one. Each node in the second rooted neighborhood has degree two. Each Node in the third rooted neighborhood cannot have degree three because ($5\choose 2$ is 10, which is the complete graph on 5 nodes, and each node having degree three would require 15 nodes.
This is not a simple Breadth First Search (BFS) algorithm. Where BFS ends is where the insight begins. First, we notice that the sizes of these neighborhoods are decreasing. That is not simply because of BFS. If we had selected a different graph, with non-decreasing neighborhoods, what we would have seen is a trivial degree-doubling node. What stopped this was that proof by contradiction assumption and the Decreasing Neighborhood Sequence Property (DNSP). When graphs do not have this property, they admit degree doubling nodes. Such trivial cases have been omitted here for simplicity, but BFS would not ignore them. Since we are in a proof by contradiction, BFS trivially producing a degree doubling node because of the structure of the graph, is not what we want. Instead we need the DNSP to constrain the neighborhood growth and require nodes to have the size of their second neighbors to be strictly less than that of their first.
Secondly, These neighborhoods act as node level identifiers, which can first help with distinguishing between different arc types. This is where the notion of interior neighbors, exterior neighbors and back arcs comes from. Further, we can use these same node level identifiers to classify triangles into six distinct types based on the endpoints of the nodes. Ultimately, we will use this classification of arcs into interior, exterior or back as one of the main resources to utilize to prove the truth of the SSNC.
A contrary but clarifying idea also began to take shape regarding the distance metric. Some authors define second out-neighbors purely in terms of reachability - a node belongs to $N^{++}(x)$ if it is reachable via a directed path of length two. There is the issue of double-counting, where a node can be a member of $N^{+}(x)$ or $N^{++}(x)$. Other authors attempt to resolve this double counting by defining second out-neighbors as the disjoint union, but doing so potentially loses track of paths that we would like to reconstruct.
If this representation could lead to a rigorous mathematical foundation, then it has the potential to both keep track of paths, and not have the problems of defining second neighbors in terms of reachability instead of shortest path. The mathematical formulation for these ’rooted’ neighborhoods is simply the shortest path distances from the chosen minimum out-degree node $v_{0}$ to each node in that neighborhood, which is the root node. This allows us to partition the graph about that node and take advantage of some of the other properties of distance, such as total order. This will be defined formally in Section 5 (Graph Level Order). As such, this will keep track of paths in our partitions, unlike previous approaches.
These initial steps were driven by intuition. These small, preliminary examples all identified degree-doubling nodes in highly predictable, structured circumstances. This does not suffice for mathematical rigor. These graphs can be deeply deceptive, as what appears obviously true or general in these limited settings may not hold under greater scrutiny. More difficult questions quickly arose like how do these graphs perform under the presence of more complex back arcs. Such arcs disrupt the interior and forward flow of arcs between rooted neighborhoods. Could we trust this visual intuition? Would these examples scale to arbitrary graphs? More importantly, how can we build a mathematical framework that generalizes this phenomena we observed? The remainder of this work is devoted to answering those questions with rigor.
Indeed, in 2020, we were confident enough in an early version of this framework to share a preliminary write-up with Dr. Paul Seymour at Princeton University. He graciously responded and pointed out a key structural issue: we had missed a case in the presence of back arcs. These are arcs directed from nodes in later neighborhoods back to earlier ones. The presence of these arcs would cause the neighborhoods to overlap, invalidating our initial attempts to treat the sets as disjoint for our inductive for additive arguments.
While our small constructed examples showed that such back arcs often quickly led to a degree doubling node, Dr. Seymour’s example was more abstract but grounded in the language of our definitions. His insight revealed that the real obstacle was not the back arcs themselves, or even the transitive triangles in which they presented themselves. We needed to prove that what we were seeing in larger, more concrete examples, would hold up in this more abstract one. This would take a more structured environment than the one we were working with. This feedback led directly to the development of the Graph Level Order data structure, which was a way to retain order, enforce layered progression, and build the foundation necessary for an algorithmic approach to the SSNC.
2.2 Set Theoretic Aspects
The next step in our analysis is the consideration of arbitrary oriented graphs. We did not want to make assumptions about node degrees or other graph properties that were too restrictive. For example we did not want to assume, at least initially, that any class of subgraphs were excluded from these oriented graphs. Instead, we wanted to consider the fundamental characteristics common to all oriented graphs. As we mentioned before, we know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree, $\delta$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
The notion of partitioning the graph into subgraphs is not introducing the NP-hard problem of set partitioning into the SSNC. The set-partition problem [13] asks if a sequence of positive numbers $S$ can be partitioned into two sets $S_{1},S_{2}$ such that the sum of the numbers in those sets are equal. That is a search for a partition. On the contrary, this algorithm begins with an anchor about the minimum out-degree node and declares that the partition be done based on that node.
By partitioning the graph, we gain the ability to make these necessary comparisons. What we see are two nodes, $u,v$ where $v\in N^{+}(u)$ (i.e., $v$ is further from $v_{0}$ than $u$). They may have shared first out-neighbors, $u$’s first out-neighbor is intersected with $v$’s second out-neighbor, $v$’s first out-neighbor is intersected with $u$’s second out-neighbor, or both second out-neighbors overlap. We thus see the following four conditions:
  • Shared First Out-Neighbors $N^{+}(u)\cap N^{+}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node, and they share common first out-neighbors. These are the shared direct out-neighbors of $u$ and $v$. This is the situation where $v$ assists with $u$ not being a degree-doubling node. This is the situation we refer to as interior neighbors.
  • $u$’s Second Neighbors vs. $v$’s First Neighbors Let $x\in N^{++}(u)\cap N^{+}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node, and $u$’s second out neighbors are common with $v$ first. These are the out-neighbors of $v$ that are reached in two steps from $u$. This gives rise to $u\to y_{1}\to x$ and $v\to x$. This is the situation where $x$ is reachable through $y_{1}$ or $v$. This illustrates what [2] defined as a Seymour diamond. Since these paths are traceable we do not lose count of the neighbors of $u$ in doing this. This is the situation we refer to as exterior neighbors.
  • $u$’s First Neighbors vs. $v$’s Second Neighbors $x\in N^{+}(u)\cap N^{++}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node and these are the out-neighbors of $u$ that are reached in two steps from $v$. This implies that we have $u\to x$ and $v\to y_{1}\to x$. Remember that $v\in N^{+}(u)$, which means that we have $u\to v$. So $x$ and $v$ have the same distance from the minimum out-degree node. This implies that the nodes $x,y_{1},z$ are all the same distance from the minimum out-degree node $v_{0}$. There are two cases to consider If $y_{1}\in N^{+}(u)$ or $y_{1}\notin N^{+}(u)$. If $y_{1}\in N^{+}(u)$ then $y_{1}\in N^{+}(u)\cap N^{+}(v)$. Similarly, if ($y_{1}\notin N^{+}(u)$, then $y_{1}\in N^{++}(u)\cap N^{+}(v)$.
  • $u$’s Second Neighbors vs $v$’s Second Neighbors $x\in N^{++}(u)\cap N^{++}(v)$:
    $u$, the parent node, is a node that is closer to $v_{0}$ than $v$, the child node and these are the common second out-neighbors of both $u$ and $v$. This would mean that they are second out-neighbors within the first out-neighborhood of $u$. Consider the node $z_{1}$. There are two possible cases. Either $z_{1}\in N^{+}(u)$ or $z_{1}\notin N^{+}(u)$. If $z_{1}\in N^{+}(u)$, then $z_{1}\in N^{+}(u)\cap N^{+}(v)$. Similarly, if $z_{1}\notin N^{+}(u)$, then $z_{1}\in N^{++}(u)\cap N^{+}(v)$.
These four intersection types allow us to track how influence and load-sharing propagate from $u$ through its neighborhood, giving a nuanced lens on why certain nodes fail to double their degree. We see from these four cases that the main two cases that need to be concentrated on are the first two, namely when there are first out-neighbors (interior neighbors), and when a parent shares second neighbors with a child’s first neighbors (exterior neighbors). The rest of this paper will delve more into these cases.
What our approach allows for is a more nuanced comparison of the out-neighborhoods through this set-theoretic approach. This allows us to not only track out-neighborhoods but also accurately evaluate the conjecture under other considerations like back arcs, transitive triangles, and Seymour diamonds, as have been discussed in previous papers.
As we noted earlier, this proof by contradiction begins at the node level. We are constructively building this house of cards with our definitions, lemmas, examples, theorems and ultimately our algorithms. Some of the first questions we will ask concern what conditions must be in place in a graph in order for a node’s degree not to double. We will begin this process with minimum degree nodes, and branch out from there. At the same time, these lemmas we are constructing are helping the house of cards to grow collectively stronger, to ultimately either produce a degree doubling node or prove that one cannot exist.
We will revisit the concept of transitive triangles, not as hindrances but as mechanisms for load sharing. These triangles allow for the neighbors of a node to support one another. This can help prevent their parent’s degrees from doubling. These shared responsibilities can prevent an immediate predecessor from being a Seymour vertex or not having their degrees double. The question then becomes: how much order is embedded in these graphs, and how frequently do these patterns emerge? Consider the example of a node $u$ with three out-neighbors $v_{1},v_{2},v_{3}$ such that we have a cycle in the subgraph $v_{1}\to v_{2}\to v_{3}\to v_{1}$. This cycle prevents node $u$’s degree from doubling. This is $u$ sharing its load with its first out-neighbors.
We will also see that node’s degree can be partitioned into three distinct components: interior neighbors, exterior neighbors, and a third called back arcs. This represents a clear distinction from standard formulations of the SSNC, where most authors consider only first and second neighbors-without a structural partitioning of the degree. We already illustrated this in Figure 1. As we will see throughout this paper, the partitioning of a node’s degree, in conjunction with the DNSP, plays a crucial role in determining how neighborhood sizes evolve as we move farther from the root node.
Complexity must be addressed when discussing global search algorithms. This raises concerns about how well they will handle larger graphs in comparison to smaller ones. This should not be a problem, as we will see in the following sections. The initial step of our algorithm involves partitioning the graph’s nodes and arcs based on their distance from the minimum out-degree node, which can be done in $O(|E|+|V|)$ time using the Breadth First Search Algorithm (BFS) which returns the levels, corresponding to the rooted neighborhoods. What we will then see is that these partitions decrease in size as we move from the minimum out-degree node. Our algorithm will simply take a path from the minimum out-degree node to the degree-doubling node.
Finally, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of Conjecture 3.1, which will be introduced in Section 3 (Graph Theory Terminology). That conjecture is looking for a node whose degree at least doubles in the square of the graph. By the equivalence of the two conjectures, this is the same as looking for a node whose second neighborhood is at least as large as its first, i.e. a node that satisfies Conjecture 1.1. A similar definition, a Seymour vertex, fulfills the conjectures—it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node. We note that we can make a degree-doubling node a noun, a property of the node having its degree double, or a verb, as in having its degree doubled. We do not have such freedom with the term Seymour vertex.
Definition 2.1.
In an oriented graph $G$, a node $v$ is a degree-doubling node (or Seymour vertex) if $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}(v)$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$ in $G$, respectively.
This structured approach, starting from a minimum out-degree node and proceeding through ordered partitions, provides a foundation for a global search algorithm designed to constructively identify a degree-doubling node in any oriented graph.
3 Graph Theory Terminology
This section is intended for readers who may not have a traditional background in graph theory. In particular, graph theory is an interdisciplinary field, finding applications in mathematics, computer science, engineering, biology, and social sciences and this proof of the SSNC has reached many of those fields. To make this paper more accessible to a broader audience, we tried to provide a concise overview of some of the key definitions and concepts. We understand that seasoned graph theorists may be familiar with much of the language used throughout, but even the concepts of oriented graphs and square graphs can span beyond a typical graph theory syllabus. Because of the nature of oriented graphs, we write $u\to v$ to indicate a directed arc from $u$ to $v$. This should help with the over-usage of parenthesis in functions and other areas of graph theory. We hope this section serves as a helpful reference point throughout the paper.
An oriented graph is a directed graph with no self-loops and no pair of vertices connected by edges in both directions—that is, at most one directed edge exists between any two vertices. The square of a graph $G$, denoted $G^{2}$, is a graph on the same vertex set where an arc exists from $u$ to $v$ if there is a path of length at most two from $u$ to $v$ in $G$. Essentially, $G^{2}$ captures all two-step connections from $u$ to $v$.
Definition 3.1.
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $u\to u$ where $u$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $u\to v$ and $v\to u$ where $u$ and $v$ are nodes in $G$.
Definition 3.2.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of arcs defined as:
$E^{2}=\{u\to v\mid u\to v\in G\text{ and }\exists w\in G\text{ such that }u\to w\to v\in G\}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$. These are the nodes that $v$ is pointing to. Similarly, $N^{++}(v)$ refers to the second out-neighbors of $v$, that is, nodes that are distance two from $v$ in $G$. We make a point to exclude the distance one node from this definition.
Definition 3.3.
The distance between nodes $u$ and $v$, denoted $\text{dist}(u,v)$, is the length of the shortest directed path from $u$ to $v$.
Definition 3.4.
The first out-neighborhood of a vertex $v\in G$ is defined as
$N^{+}(v)=\{w\in G\mid v\to w\in G\}$
Definition 3.5.
The second out-neighborhood of a vertex $v\in G$ is defined as
$N^{++}(v)=\{u\in G\mid\exists w\in G\text{ such that }v\to w\to u,\text{ and }u\notin N^{+}(v)\}$
Definition 3.6.
Let $G=(V,E)$. Let $S\subseteq V$ be a subset of the vertices of $G$. Then the induced subgraph $G[S]$ is the graph whose vertex set is $S$ and whose edge set consists of all the edges in $E$ that have both endpoints in $S$.
For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $\text{dist}(u,v)$ is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 3.1.
[…]
Figure 2: A five-vertex oriented graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices $A$, $B$, and $C$.
Conjecture 3.1.
(Square version of Seymour’s Second Neighborhood Conjecture). For every oriented graph $G$, there exists a vertex $v$ such that:
$d^{+}_{G^{2}}(v)\geq 2\cdot d^{+}_{G}(v)$
Conjecture 1.1 relates the degree doubling condition to a vertex’s neighborhood growth, while Conjecture 3.1 frames the same idea in terms of the squared graph’s structure. Since each arc in $G^{2}$ corresponds to a path of length two in $G$, the two conjectures are logically equivalent. For clarity and practicality, much of our work focuses initially on Conjecture 1.1, as neighborhood-based reasoning provides more intuitive and accessible insights. (See Conjecture 1.1 in Section 1) However, later proofs will also directly invoke Conjecture 3.1.
Example 3.2.
12345
Figure 3: A five-vertex oriented graph showing original edges with the full edges and the square cycles with the dashed edges. Every node satisfies both versions of the conjecture since they all have out-degree of 1.
We can illustrate the equivalence between Conjectures 1.1 and 3.1 by showing that the same nodes satisfy the degree-doubling condition in both cases. To do so we will consider a graph $G$ that consists of a cycle on $n$ nodes and $n$ arcs, represented as $v_{0}\to v_{2}\to v_{3}\to\dots\to v_{n-1}\to v_{0}$. We can see this through Example 3.2 with a graph on five nodes.
First, we will consider Conjecture 1.1. Let $v_{i}$ be a node on this cycle. Then we know its first out-neighbors $N^{+}(v_{i})=\{v_{i+1}\}$ (Indices are modulo $n$). Its set of second out-neighbors is $N^{++}(v_{i})=\{v_{i+2}\}$. Thus, we have that $|N^{++}(v_{i})|=1$ and $|N^{+}(v_{i})|=1$. This means that there are at least as many second out-neighbors as first out-neighbors, and every node $v_{i}$ in the cycle satisfies the requirements of the conjecture $|N^{++}(v_{i})|\geq|N^{+}(v_{i})|$.
Next, we consider the same cycle in Conjecture 3.1. For every node $v_{j}$ in the directed cycle, the out-degree in $G$ is $d^{+}_{G}(v_{i})=1$. In the squared graph $G^{2}$, there are arcs $v_{i}\to v_{i+1}$ (modulo n) as well as $v_{i}\to v_{i+2}$, so the out-degree is $d^{+}_{G^{2}}(v_{i})=2$. Thus we have $d^{+}_{G^{2}}(v_{i})\geq 2\cdot d^{+}_{G}(v_{i})$ in this case.
The directed cycle is just one example where a graph satisfying Conjecture 1.1 also inherently satisfies Conjecture 3.1. In this specific example, the set of nodes satisfying both conjectures is the entire vertex set of the cycle. This highlights how the fundamental requirement of having a sufficient number of second out-neighbors is central to both perspectives of the conjecture.
Therefore, while we will utilize both conjectures, our initial focus will be on Conjecture 1.1 due to its compatibility with neighborhood-based reasoning and the intuitive insights it offers. Nonetheless, we will also directly employ Conjecture 3.1 in several of our proofs.
For readers seeking a more formal or comprehensive introduction to graph theory, we recommend standard references such as Reinhard Diestel’s Graph Theory [10] provides a thorough and widely used introduction to the standard foundations of the field. Graphs, Networks and Algorithms [19] by Dieter Jungnickel approaches the subject from a computer science perspective.
4 Initial Lemmas
What follows is a set of lemmas that seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. Unless otherwise noted, we assume that $v_{0}$ is a node with the minimum out-degree in our oriented graph $G$. The work in this section agrees with what was already done by [20] and is not intended as original content. We provide it here to guide readers into our proof by contradiction methods. As we stated in the introduction, we begin with a single node, asking what requirements need to be in place for that node’s degree not to double. If we are able to successfully achieve a single node whose degree does not double, and list the accompanying requirements, we would like to continue to move to other nodes in the graph.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
Figure 4: This is an example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$. There is a neighbor of degree 0 in the neighbor induced subgraph. This scenario shows how even with a minimum degree node with out-degree 3, the presence of a neighbor with no internal connectivity guarantees an unavoidable increase in second neighbors.
Lemma 4.1.
Minimum Out-Degree $<$ 3 Suppose that the minimum out-degree node $v_{0}$ in the oriented graph $G$ has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. In this case $v_{0}$ has no out-neighbors, i.e, $|N^{+}(v_{0})|=0$. Since $v_{0}$ has no out-neighbors, it also has no second out-neighbors. This means that $|N^{++}(v_{0})|=0$. This will imply that $|N^{++}(v_{0})|=|N^{+}(v_{0})|$, which will make $v_{0}$ a degree-doubling node.
Case 2: Assume that $d^{+}(v_{0})=1$. Here, $v_{0}$ has a unique out-neighbor. Let $v_{1}$ be that unique out-neighbor of $v_{0}$. We know that $|N^{+}(v_{0})|=1$. Because $v_{0}$ is a minimum out-degree node, the out-degree of $v_{1}$ must be at least 1. This means that $v_{1}$ must have an out-neighbor. This neighbor cannot be $v_{0}$ since this graph is oriented and $v_{0}$ already has an arc to $v_{1}$. This means that it must be some node $v_{2}$, which will be a second out-neighbor of $v_{0}$. So we know that $|N^{++}(v_{0})|\geq 1$. Thus we have that $|N^{++}(v_{0})|\geq|N^{+}(v_{0})|$ and $v_{0}$ is a degree-doubling node.
Case 3: Assume that $d^{+}(v_{0})=2$. Let’s call the out-neighbors of $v_{0}$, $v_{1}$ and $v_{2}$, where $|N^{+}(v_{0})|=2$. We know that $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$ since $v_{0}$ is a minimum out-degree node. Consider the induced subgraph $G[v_{1},v_{2}]$. At most one arc can exist in an oriented graph of size 2 because oriented graphs do not allow for symmetric arcs. So we know that there is at most one arc between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that $v_{1}$ or $v_{2}$ will have at least two arcs outside of $G[v_{1},v_{2}]$. Neither $v_{1}$ or $v_{2}$ can relate to $v_{0}$ since $G$ is an oriented graph and $v_{0}\to v_{1}\in G$ and $v_{0}\to v_{2}\in G$. We can define the nodes $v_{3}$ and $v_{4}$ as the two out-neighbors of $v_{1}$ not in ${v_{0},v_{1},v_{2}}$. This means that $v_{0}$ will have $v_{3}$ and $v_{4}$ as second neighbors. Hence, $v_{0}$ will have at least two second out-neighbors—$v_{3}$ and $v_{4}$. Thus $|N^{++}(v_{0})|\geq 2$ and $|N^{++}(v_{0})|\geq|N^{+}(v_{0})|$, and $v_{0}$ is a degree-doubling node.
Therefore, in all cases, $v_{0}$ is found to be a degree-doubling node when its minimum degree is less than 3. ∎
Lemma 4.1 (Minimum Out-Degree $<$ 3) begins to tell us about the necessary conditions for degree doubling nodes in oriented graphs. What we see from this first lemma is that we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum out-degree node has very few out-neighbors, not enough to support the load balancing that we will see is necessary to prevent the degree doubling of nodes.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 in the Neighbor Induced Subgraph
Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3. Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0 in its neighbor induced subgraph. Then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{1},v_{2},v_{3}$ be the out-neighbors of $v_{0}$. We see then that $|N^{+}(v_{0})|=3$. Consider the neighbor induced subgraph $G[v_{1},v_{2},v_{3}]$.
Assume that $v_{1}$, which is a neighbor of $v_{0}$ has out-degree 0 in $G[v_{1},v_{2},v_{3}]$. This means that $v_{1}\to v_{2}\notin G[v_{1},v_{2},v_{3}]$ and $v_{1}\to v_{3}\notin G[v_{1},v_{2},v_{3}]$. Since $v_{0}$ is a minimum out-degree node with out-degree 3, we know that $v_{1}$ must have out-degree at least 3. So $v_{1}$ must connect to at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. Since $v_{0}\to v_{1}\in G$, we know that $v_{1}\to v_{0}\notin G$ by the oriented property of $G$. This means that $v_{1}$ must connect to at least three other nodes $v_{4},v_{5},v_{6}$. So we have $|N^{+}_{G}(v_{1})|\geq 3$. The nodes $v_{4},v_{5},v_{6}$ are not in $G[v_{1},v_{2},v_{3}]$ and are not equal to $v_{0}$, so they are outside $G[v_{0},v_{1},v_{2},v_{3}]$.
This means that these first out-neighbors of $v_{1}$ are second out-neighbors of $v_{0}$, so we have that $|N^{++}_{G}(v_{0})|\geq 3$. We can simplify this and say that $|N^{++}(v_{0})|\geq|N^{+}(v_{0})|$. This means that $v_{0}$ is a degree-doubling node. Hence, $v_{0}$ must be a degree-doubling node, contradicting the assumption that such a node could exist in a minimal counterexample. ∎
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0 in the Neighbor Induced Subgraph) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first out-neighbors must be $v_{0}$’s second out-neighbors. By assumption, $v_{1}$ must have at least as many first out-neighbors as $v_{0}$, and since these are not first out-neighbors of $v_{0}$, they will be second out-neighbors of $v_{0}$, causing $v_{0}$’s degree to double. This scenario shows how even with degree 3, the presence of a neighbor with no internal connectivity guarantees an unavoidable increase in second out-neighbors.
Example 4.2.
$v_{0}$$v_{2}$$v_{1}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$
Figure 5: Example illustrating out-neighborhood partitioning and how a node in $G$ has its degree doubled in $G^{2}$ when a minimum-out-degree node has all its neighbors have degree 1.
Lemma 4.3.
Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph Assume that in an oriented graph $G$, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of $v_{0}$’s out-neighbors, every neighbor has an out-degree of 1. Then the out-degree of $v_{0}$ or one of its neighbors will be a degree-doubling node.
Proof.
Assume $d^{+}(v_{0})=3$. Let $v_{1},v_{2},v_{3}$ be the out-neighbors of $v_{0}$. Then $|N^{+}(v_{0})|=3$. We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, $v_{0}$ will have a larger second out-neighborhood than its first out-neighborhood, making it a degree-doubling node. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$. $|N^{+}(v_{i})|=3$.
Similarly, by Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), We see that if any of the three first out-neighbors of $v_{0}$ does not have any first out-neighbors of $v_{0}$ as an out-neighbor, then $v_{0}$ has its degree doubled.
To prevent $v_{0}$ from being a degree-doubling node, each of the first out-neighbors of $v_{0}$ needs to have another first out-neighbor of $v_{0}$ as an out-neighbor. What this means is that for those nodes that are first out-neighbors of $v_{0}$, we have $|N^{+}_{G-G[v_{1},v_{2},v_{3}]}(v_{i})|=2$.
This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one arc between $v_{4}$ and $v_{5}$. We will say that $v_{4}$ has three additional out-neighbors, $v_{6},v_{7}$, and $v_{8}$, with $v_{4}$ having degree zero in $G[v_{4},v_{5}]$. Then $|N^{+}(v_{4})|=3$. These three neighbors of $v_{4}$ are second out-neighbors of $v_{1}$, giving us $|N^{++}(v_{1})|=3$. Thus we have the following equality, $|N^{++}(v_{1})|=|N^{+}(v_{1})|=3$, making $v_{1}$ a degree-doubling node.
Hence, $v_{1}$ must be a degree-doubling node, contradicting the assumption that such a node could exist in a minimal counterexample. ∎
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph) hints at a strategy of examining the out-neighborhoods of neighbors, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling for it to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
We did not require the right starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimum out-degree node due to its ability to split the graph’s nodes. We can see an illustration of Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph) in Example 4.2.
These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
These initial lemmas establish that any graph serving as a minimal counterexample to the SSNC must possess a minimum out-degree of at least 3. Furthermore, even with this minimum degree, specific local structures within the out-neighborhood of a minimum out-degree node can guarantee the existence of a degree-doubling node, suggesting constraints on the architecture of potential counterexamples.
5 Graph Level Order
In the world of computer science, data structures are a fundamental tool for organizing information. Efficient data structures have helped not only provide much of how the real world works today. Not only that, but many of the major open problems were helped with data structures. The development of balanced trees (AVL trees, Red-Black trees, and B-trees) helped with efficient searching and information retrieval. Graph data structures have been crucial for algorithms like max flow/min cut problems and network optimization problems. Stacks, queues and parse trees helped with compiler design and programming languages. Data structures have also been front and center in the world of open problems.
We have begun a bottom up approach to the SSNC, starting with a minimum degree node. What we are going to do is continue this approach, culminating in a data structure that will order the nodes by distance from that minimum degree node. Conceptually, this is how Example 4.2 was drawn, but we will make this a part of our formal notation. By doing this, we will be able to take advantage of other properties of this data structure and its organization of the nodes within it.
The Example 4.2 is not the first time we have seen the minimum out-degree node not having its degree double. Fisher also provided an example in [11] where the minimum out-degree node does not have its degree double. In Example 4.2, this node happens to be a first out-neighbor of the minimum out-degree node. There is no lasting guarantee that this will always be the case. There is no reason why future examples would have to stop in first out-neighbors of minimum out-degree nodes. This process could continue through a sequence of out-neighbors of the minimum out-degree node. No proof has been given that this process has to stop anywhere just yet.
Up until now, we have been using the terminology of induced subgraphs to represent adjacent out-neighborhoods. We have just begun to speak of why this might not be sufficient for what we have planned. In Example 4.2, it was not difficult to speak of the single induced subgraph and the next induced subgraph using $G[v_{1},v_{2},v_{3}]$ and $G-[v_{1},v_{2},v_{3}]$, respectively. If this degree-doubling node had been a further distance from $v_{0}$, expressing it through such an induced subgraph representation would not have been so easy. As the distance between the degree-doubling node and the minimum out-degree node $v_{0}$ is increased, the number of induced subgraphs begins to become longer, and this notation becomes challenging. What might begin to happen is that notation becomes confusing or intent begins to get in the way over innovation, so time spent discussing ideas for a proof is instead spent trying to understand parentheses to ensure that we are all in the correct ’induced subgroup’. Instead, we plan to introduce a new terminology that covers our thought process, and ’keep the main thing the main thing’.
5.1 Definitions
In this section, we develop a new data structure for oriented graphs, Graph Level Order. This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as rooted neighborhoods. Additionally, we shall formally define interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the path-finding technique to help solve the SSNC.
Definition 5.1.
Given an oriented graph $G$ and a minimum out-degree node $v_{0}$, a rooted neighborhood $R_{i}$ of distance $i$ is the subgraph of $G$ induced by the set of all nodes at distance exactly $i$ from $v_{0}$. Formally, let
$N_{i}^{+}(v_{0})=\{v\in G\mid\text{dist}(v_{0},v)=i\}$
Then, $R_{i}=G[N_{i}^{+}(v0)]$.
Figure 6: Illustration of a rooted neighborhood. Rooted neighborhoods group nodes based on their distance from the minimum out-degree node $v_{0}$, with no assumptions on the sizes of these rooted neighborhoods.
Similar constructions have appeared in prior work. For instance, Daamouch employed induced subgraphs in his analysis of $m$-free graphs [6]. However, the focus here is more sharply placed on the hierarchical structure of rooted neighborhoods (see Example 6), which plays a central role in our proofs. This is also closely related to the notion of the $i$-th rooted neighborhood introduced independently by Botler et al. [1], who defined neighborhoods in terms of shortest paths. Their work focused on the global behavior of (pseudo)random graph orientations. While their results are powerful and address large-scale properties, our aim is to explore the fine-grained, local behavior of these rooted neighborhoods. We propose that understanding these local node interactions will give rise to clearer global insights of the conjecture.
We investigate oriented graphs from two perspectives: through the lens of rooted neighborhood layers and through the position of a node relative to a fixed minimum out-degree vertex $v_{0}$. Each node belongs to a rooted neighborhood at a specific distance from $v_{0}$ and is either before, in, or after a given rooted neighborhood when viewed through this layering.
One key difference in our approach lies in the utilization of set theory to extend some of the previous concepts of graph theory. We believe there is a rich structure within these rooted neighborhoods—a structure that has not yet been fully explored—and that this perspective allows us to ask and answer fundamentally different questions than previous work. The notions of induced subgraphs and disjoint subgraphs, as defined previously, do not have the anchoring property of the minimum out-degree node. We could have simply stuck with standard notation, but attempting to use induced subgraphs while referencing a minimum out-degree node gets readers trapped in notation while trying to convey an important concept.
These rooted neighborhoods are anchored at a minimum out-degree node, and their importance is revealed by the element $i$, which identifies the distance between that rooted neighborhood and this minimum out-degree node. This allows us to not only partition the nodes, which we have already done by the rooted neighborhoods, but now we can keep track of entire paths from $v_{0}$ to every node in every $u_{i}\in R_{i}$. These nodes within the rooted neighborhoods will have hierarchical relationships. That is, relationships with nodes in the rooted neighborhoods closer to $v_{0}$, relationships with nodes in their own rooted neighborhood, and relationships with nodes in rooted neighborhoods further from $v_{0}$. These relationships are the backbone of the Graph Level Order. Within this section, we will show that this is not just a labeling. There are important properties that are revealed within these rooted neighborhoods being at distance $i$ from the minimum out-degree node.
At the heart of these rooted neighborhoods are parent-child relationships. A node $u_{i}$ at level $R_{i}$ has out-neighbors at level $R_{i+1}$. These out-neighbors at the next level $v_{1},v_{2},\ldots,v_{n}$, which are the children of $u_{i}$, have a sibling relationship to each other. In this proof by contradiction, the goal of the parent at each level $G[N^{i+}]$ is to prevent their degree from doubling. Demonstrating how their children, through a process of interlocking, help share that load will show this is impossible. The next section formalizes this, detailing the graph-theoretic specifics of the interlocking and how it prevents degree doubling. First, however, the meaning of ”parent,” ”child,” ”sibling,” and ”the load” in this context will be clarified.
Definition 5.2.
Transitive Triangle Let $x\to y,x\to u,y\to u$. Then $x$, $y$, and $u$ form a transitive triangle.
Definition 5.3.
Parent and Child Let $u_{i}$ be a node in the rooted neighborhood $R_{i}$ for some $i\geq 0$. A child of $u_{i}$ is a node $v_{i+1}\in R_{i+1}]$ such that $u_{i}\to v_{i+1}$ (we also say that $u_{i}$ is the parent of $v_{i+1}$).
Definition 5.4.
Interior Neighbor and Interior Degree Let $u_{i}\in R_{i}$ be a parent node with children $v_{1},v_{2}\in R_{i+1}$. We define the interior neighbors of $v_{1}$ with respect to $u_{i}$ as those nodes $z\in G$ such that both $u_{i}\to z$ and $v_{1}\to z$. That is, nodes that are common out-neighbors of both $u_{i}$ and $v_{1}$, forming transitive triangles.
$\text{int}(u_{i},v_{1}):=N^{+}(u_{i})\cap N^{+}(v_{1})$
The interior degree of $v_{1}$ with respect to $u_{i}$ is defined as
$|\text{int}(u_{i},v_{1})|$
Definition 5.5.
Siblings Let $u_{i}\in R_{i}$ be the parent of $v,w\in R_{i+1}$. Then $v$ and $w$ are said to be siblings.
Now that we have introduced neighbors, we can talk about parent-child relationships between rooted neighborhoods. This happens when a node in one rooted neighborhood $R_{i}$ has one or more out-neighbors in the next rooted neighborhood $R_{i+1}$. What this establishes is the possibility to discuss relationships between these common out-neighbors of the parent and children. For example, in Example 4.1, we saw that node $v_{0}$ was a parent of nodes $v_{1},v_{2}$, and $v_{3}$, but only had common out-neighbors in rooted neighborhood $R_{1}$ with nodes $v_{2}$ and $v_{3}$. The node $v_{2}$ is called an interior neighbor of $v_{3}$ in this example.
With the concept of transitive triangles now defined in the context of parent, child, and child-neighbor relationships, we can begin to examine the role of children in supporting their parents. To approach the problem from a proof by contradiction angle, we consider a start node that we would like to avoid having its degree double. Rather than describing this in strict mathematical terms, we can frame it metaphorically: the node (or parent) does not want to carry a heavy burden. We think of this as something that could potentially overwhelm it, like becoming a degree-doubling node.
In this context, the children play a critical role. These child nodes can volunteer to share in the responsibility by reducing the parent’s load. This collaborative dynamic forms the foundation of what we refer to as load balancing. When a parent node is supported by its children through transitive relationships—effectively redistributing influence or responsibility—this can prevent the parent from becoming a bottleneck or singular point of stress in the graph.
In approaching the SSNC conjecture from a constructive and contrapositive perspective, transitive triangles emerged almost immediately within the first neighborhood of many vertices. These structures provided a natural way to understand interior support via shared out-neighbors—what we now call interior neighbors. However, the SSNC conjecture is not only a statement about the first neighborhood $N^{+}(v)$ of a vertex $v$, but also it joins them with second neighborhoods $N^{++}(v)$. This requirement prompted a deeper investigation of what patterns should be emerging inside these graphs. If we are seeing transitive triangles represented among the first neighbors, what does this say about the second neighbors?
Because these second neighbors are akin to the counterparts to the first neighbors, they should have relationships to those same parent $u$ and child $v$ nodes in the transitive triangle, but from a different angle. They are second neighbors of the parent node $u$, so the conjecture itself will bound the size of these elements. Likewise, these exterior neighbors will be related to the child node $v$ because there must be a path from $u\to v\to w$ for any exterior neighbor $w$.
While the interior neighbors portrays a concept similar to transitive triangles, the notion of exterior neighbors is entirely new. The core of the idea is that once we have these rooted neighborhoods, a node’s degree can be split into interior and exterior degrees. Within those two degrees, we expect two different types of behaviors. One type of behavior, interior degree, we expect to be able to show is easy, predictable, and tractable behavior. This is because these interior neighbors, relative to the parent nodes, will cause cycles among the interior neighbors to prevent that parent from becoming a degree doubling node. Exterior degree, on the other hand, will be a more strenuous effort. To understand how exterior neighbors impact the conjecture, we need to see how they interact with the Graph Level Order and the Decreasing Neighborhood Sequence Property (DNSP).
Definition 5.6.
Let $u_{i}\in R_{i}$ be a parent of a node $v_{i+1}\in R_{i+1}$. The exterior neighbors of $v_{i+1}$ with respect to $u_{i}$ are nodes $z$ such that $z$ is a second out-neighbor of $u_{i}$ and a first out-neighbor of $v_{i+1}$, i.e., $z\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$. This implies that there exists a path $u_{i}\to w\to z$, and an arc $v_{i+1}\to z$ exists, but $u_{i}\to z\notin G$. Unlike the interior neighbors, exterior neighbors are neighbors of the child that are not shared by the parent.
The exterior degree of $v_{i+1}$ with respect to $u_{i}$ is defined as
$|\text{ext}(u_{i},v_{i+1})|.$
Exterior neighbors reveal how a rooted neighborhood interacts with adjacent rooted neighborhoods. This interaction is critical for the SSNC, as the nodes within the current rooted neighborhood depend on the connections to the next rooted neighborhood. We return to the Example 4.1 to see how the node $v_{1}$ treated its parent node $v_{0}$. In that situation, $v_{0}$ and $v_{1}$ shared no common neighbors; thus, $\text{int}(v_{0},v_{1})=\emptyset$. As a result, we have that $\text{ext}(v_{0},v_{1})=d^{+}(v_{1})=\{v_{4},v_{5},v_{6}\}$, i.e. $|d^{+}(v_{0})|=|\text{ext}(v_{0},v_{1})|=3$. Contrast that with Example 4.3, where we saw that $\text{int}(v_{0},v_{1})=\{v_{2}\}$, $\text{int}(v_{0},v_{2})=\{v_{3}\}$, $\text{int}(v_{0},v_{3})=\{v_{1}\}$. Similarly, in this example, the nodes $v_{1},v_{2},v_{3}$ would be siblings.
Notice that interior out-neighbors are only defined for rooted neighborhoods 1 and greater. This is because the minimum out-degree node $v_{0}$ is the anchor. It has no neighbors in $R_{0}$. As such, all its arcs are exterior arcs.
Remark 5.1.
For the minimum out-degree node $v_{0}$, all its out-neighbors are exterior neighbors because $R_{0}=\{v_{0}\}$ and there are no earlier neighborhoods $R_{-1}$. Thus,
$\text{int}(v_{0},v)=\emptyset\hskip 28.45274pt\forall v\in R_{1}$
The concept of interior neighbors, as we define it, represents a specific and crucial type of transitive triangle that arises within our Graph Level Order. While Brantner et al. [2] were instrumental in highlighting the general significance of transitive triangles—which can also arise from back arcs and present hindrances to the conjecture and exterior neighbors, which do not present a problem to the conjecture—and introduced the related idea of shared first out-neighbors, our formulation introduces a critical distinction. Our definition of interior neighbors specifically focuses on the directed arcs between the children of a common parent within the Graph Level Order and how they relate to that parent. Furthermore, Brantner et al.’s work did not extend to a parallel formulation for exterior neighbors.
We postulate that, much like the necessity of studying both primal and dual programs to achieve a comprehensive understanding in linear programming [24], a complete resolution of the SSNC requires the dual analysis of both interior and exterior neighbors. As we will demonstrate throughout this paper, the properties of these interior and exterior neighbors—which are fundamentally defined by their relationship to a common parent within the Graph Level Order—will provide us with crucial bounds on the actions and limitations of that parent node and its implications for the conjecture.
Exterior neighbors do relate to what Brantner et at. [2] referred to as Seymour diamonds. These are subgraphs where a node $x$ connects to a node $z$ through two different paths $y_{1}$ and $y_{2}$, $x\to y_{1}\to z$ and $x\to y_{2}\to z$. Seymour diamonds presented a problem for the SSNC because the dual paths to the node $z$ meant that it could be improperly counted as two second neighbors of the node $x$.
This concept is related to exterior neighbors because in the Graph Level Order, the node $x$ would be a parent node, with the children $y_{1}$ and $y_{2}$ who are siblings. These children can then serve as an exterior neighbor to $z$, (i.e $\text{ext}(x,y_{1})=z$ and $\text{ext}(x,y_{2})=z$. This lets us know that $z$ is a second neighbor of $x$, but with trackers through the nodes $y_{1}$ and $y_{2}$, which resolves the issue of over-counting $z$.
Figure 7: Illustration of interior and exterior neighbors. Node 00 is the parent node. The nodes $1\to 2\to 3\to 1$ forming a cycle are the first neighbors (or children) of node 00. The interior neighbors are always distance one from the parent node. The nodes $\{4,5\}$ are the exterior neighbors from any arc that parent node 00 and a node in $R_{1}$. The exterior neighbors are always second neighbors of the parent node.
Definition 5.7.
Let $v_{0}$ be a minimum out-degree node. Suppose that $x$ is a node in the rooted neighborhood $R_{i}$. A back arc is defined as an arc $x\to y$ such that $y\in N^{+}(x)$ and $y\in R_{j}$, where $j<i$.
Figure 8: Illustration of a node in rooted neighborhood $R_{3}$ with a back arc to the rooted neighborhood $R_{1}$.
Lemma 5.1.
Partition of Node’s Degree For any $v\in R_{i+1}$ with parent $u\in R_{i}$, the out-neighbors of $v$ can be partitioned as:
$N^{+}(v)=\text{int}(u,v)\cup\text{ext}(u,v)\cup\text{back}(v)$
where the three sets are pairwise disjoint.
Proof.
Let $v\in R_{i+1}$ be a node with parent $u\in R_{i}$. We have defined the following three sets for any out-neighbor $w\in N^{+}(v)$:
  • $\text{back}(v)=\{w\in N^{+}(v)\mid\text{dist}(v_{0},w)<\text{dist}(v_{0},v)\}$
  • $\text{int}(u,v)=\{w\in N^{+}(v)\mid\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land u\to w\in G\}$
  • $\text{ext}(u,v)=\{w\in N^{+}(v)\mid(\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land u\to w\notin G)\lor(\text{dist}(v_{0},w)>\text{dist}(v_{0},v))\}$
We now show that these three sets form a partition of $N^{+}(v)$. Let $w\in N^{+}(v)$. The distance from $v_{0}$ to $w$, $\text{dist}(v_{0},w)<\text{dist}(v_{0},v)$, $\text{dist}(v_{0},w)=\text{dist}(v_{0},v)$, or $\text{dist}(v_{0},w)>\text{dist}(v_{0},v)$. This gives us three exhaustive and mutually exclusive cases for each $w\in N^{+}(v)$.
Case 1: $\text{dist}(v_{0},w)<\text{dist}(v_{0},v)$ By definition, $w\in\text{back}(v)$. Such an arc $v\to w$ is a back arc in the rooted neighborhood structure. Since its distance from $v_{0}$ is less than $\text{dist}(v_{0},v)$, it cannot satisfy the conditions for $\text{int}(u,v)$ or $\text{ext}(u,v)$.
Case 2: $\text{dist}(v_{0},w)=\text{dist}(v_{0},v)$ In this case, $w$ is in the same rooted neighborhood $R_{i+1}$ as $v$. We consider the relationship between $u$ (the parent of $v$ in $R_{i}$) and $w$:
  • If $u\to w\in G$. This means $w$ is an out-neighbor of $v$ within the same layer, and also a direct out-neighbor of $v$’s parent $u$. So $w\in\text{int}(u,v)$
  • If $u\to w\notin G$. This means $w$ is an out-neighbor of $v$ within the same layer, but not a direct out-neighbor of $u$. So $w\in\text{ext}(u,v)$
These two sub-cases are mutually exclusive, so $w$ is uniquely assigned.
Case 3: $\text{dist}(v_{0},w)>\text{dist}(v_{0},v)$ Since $v\to w\in G$ is an arc, and we are in a shortest-path graph from $v_{0}$, the only possible distance relationship for $w$ to be greater than $v$’s distance is $\text{dist}(v_{0},w)=\text{dist}(v_{0},v)+1$. This implies $w\in R_{i+2}$. Since $u\in R_{i}$ and $w\in R_{i+2}$, there cannot be a direct arc $u\to w\in G$ in a shortest-path graph (as this would imply $\text{dist}(v_{0},w)=\text{dist}(v_{0},u)+1=i+1$, contradicting $\text{dist}(v_{0},w)=i+2)$. Thus, $u\to w\notin G$. By definition, $w\in\text{ext}(u,v)$.
In every exhaustive case, we were able to show that $w$ belongs to exactly one of the sets $\text{int}(u,v)$, $\text{ext}(u,v)$, or $\text{back}(v)$. Therefore, these three sets form a pairwise disjoint partition of $N^{+}(v)$. ∎
Lemma 5.1 (Partition of Node’s Degree) is showing that these interior neighbors, exterior neighbors, and back arcs are fundamentally linked in a much stronger way that first and second neighbors are not. That is, they form a partition of a node’s degree. With that, any question about the redundancy of these definitions should be erased. Secondly, we need to understand that this lemma is dissecting the child node’s degree of an arc $u\to v$, where the node $u$ belongs to an earlier neighborhood. The node partition lemma shows that the node $v$ can now have its degree partitioned into interior neighbors, exterior neighbors, and back arcs. The node $u$ can have its degree partitioned as well, but it would need to be based on its own parents from previous rooted neighborhoods.
We have now laid the groundwork to discuss our approach. Suppose we have a node $u$ whose degree we want to control. We have a node $v$ that is an out-neighbor of $u$. Now that we have defined the three ways we can divide degree-interior, exterior, and back arcs-we see how a node’s degree will be partitioned among those nodes. Each of those sets is disjoint because the rooted neighborhoods, the nodes are being sent to, are disjoint. Further, we will spend a lot of time excluding ourselves from the set of back arcs. We will divide $v$’s neighbors into the two remaining sets: those that are also neighbors of $u$ and those that are not neighbors of $u$. Thus we will focus our energy on dividing a node $v$’s out-neighbors that come from a parent $u$ into interior out-neighbors within the same rooted neighborhood and exterior out-neighbors in the next rooted neighborhood.
Because of the disjoint nature of our rooted neighborhoods, the interior and exterior neighbors are also disjoint. This is because while the interior neighbors deal with the first neighbor of a parent, the exterior neighbor deals with the second neighbor of that same parent. These neighborhoods are defined to be disjoint from one another when there are no back arcs, causing the interior and exterior neighborhoods to be disjoint. We can also say that for a given node $v$ and a parent $u$, in the absence of back arcs, that node’s entire out-degree can be represented by its interior and exterior arcs through that parent.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a rooted neighborhood’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in section 7 (Back Arcs).
Next, we represent a data structure that is the combination of the ideas we have presented. This data structure should not be a surprise: It is just stating facts that we know about distance being a total order, and what it does to rooted neighborhoods, and nodes within those rooted neighborhoods. That is, by having a total order, we have a universal ordering of both of these claims. Secondly, as the SSNC asks for another metric to compare nodes (like degree or lexicographic order), this data structure allows us to compare nodes within the rooted neighborhoods by their degrees as well. We also see that there are definitions of interior and exterior out-neighbors inherited from the graph itself.
Definition 5.8.
A Graph Level Order (GLOVER) on a directed graph $G=(V,E)$ can be defined as follows:
  1. 1.
    Leveled Rooted Neighborhood Structure: The vertices of $V$ with minimum out-degree node $v_{0}$ are partitioned into levels of Rooted Neighborhoods $R_{1},R_{2},\ldots,R_{n}$, where $R_{i}=\{v\in G:\text{dist}(v_{0},v)=i\}$, and $\text{dist}(v_{0},v)$ is the shortest path from $v_{0}$ to $v$.
  2. 2.
    Universal Rooted Neighborhood Order: The rooted neighborhoods are totally ordered such that $R_{i}<R_{j}$ if and only if $i<j$.
  3. 3.
    Comparability Within Rooted Neighborhoods: For any two vertices $u,v\in R_{i}$, their order is determined based on a specific metric (e.g., degree).
  4. 4.
    Universal Vertex Order: For any two vertices $u\in R_{i}$ and $v\in R_{j}$ with $i<j$, $u$ is considered less than $v$.
  5. 5.
    Interior and Exterior Out-neighbors: For a node $u\in R_{i}$ and $v\in R_{i+1}$, where $u$ is the parent of $v$
    • The interior neighbors of $u$ and $v$ are defined by the set $\text{int}(u,v)$.
    • The exterior neighbors of $u$ and $v$ are defined by the set $\text{ext}(u,v)$.
The Graph Level Order (GLOVER) is a novel and essential data structure at the heart of our proof of the SSNC. Before this data structure even begins with a BFS-style partition, we have to ensure that, under the proof by contradiction assumption, these neighborhood sizes exhibit a Decreasing Neighborhood Sequence Property (DNSP). That is, each node must have a second neighborhood that is strictly smaller than its first neighborhood. It is only then that the data structure can proceed to partitioning the graph by Breadth-First Search (BFS) distance from a minimum out-degree node, and further organizing vertices within each partition by a second metric, such as degree. This is not simply an artifact of BFS. The absence of this critical DNSP step could directly lead to a degree-doubling node, and harm the construction of proof by contradiction.
Fundamentally, the Graph Level Order is far more than a simple sorting of nodes by BFS distance. Rather, in order to construct this data structure, we need both an understanding of the BFS algorithm, the DNSP, and how they impact the overall SSNC. The BFS algorithm simply determines a graph’s shortest path layers. The Graph Level Order gives those layers meaning, whether those meanings are trivially a degree doubling node, or the ability to dissect a node’s degree into core components: interior neighbors, exterior neighbors and back arcs, which was not possible with first and second neighbors. It gives nodes identifiers based on their distance from the root node. These identifiers allow us to dissect transitive triangles into six distinct types, so that instead of treating them uniformly and seeking graphs without transitive triangles, we can plan for each type. This data structure defines exterior neighbors so that we can now have trackers on Seymour diamonds, and they are no longer double counted second neighbors. It also defines rooted neighborhoods, and we can show that these decrease in size, which is the fundamental result we need to show that back arcs are not a problem for the conjecture.
This methodology, which extends foundational algorithmic concepts, is consistent with established practice in graph theory. Many previous authors have considered the BFS algorithm as a first step in research. Consider the Hopcroft-Karp algorithm for maximum matching [4]. This is very efficient and one of the most widely used algorithms for finding maximum matching today. The algorithm begins by conducting a BFS on the graph. Initially, algorithms would only add one path at a time, leading to a complexity of $O(m\cdot n)$. This algorithm came along and was built on top of BFS. Instead of doing one path at a time, it found all shortest paths in parallel and did them in one phase, cutting the run time to $O(\sqrt{n}\cdot m)$.
The Graph Level Order compares directly to this, as a central example of this is the distinction between interior, exterior neighbors, and back arcs. As we saw in Lemma 5.1 (Partition of Node’s Degree), we are first able to split a node’s degree equation into interior degree, exterior degree and back arc degree. We will then prove, using abstract algebra, that the interior degree of a node will double if it needs to map to $i$ other nodes. Next, we will prove that the size of the exterior neighbor sets will decrease along any path from the root node $v_{0}$ to any node on that path, as long as there are no back arcs. Then we will generalize the load balancing concept we have been talking about to show that every node $u_{i}\in R_{i}$ must have interior degree $i$.
This will leave us ready to develop a traversal algorithm through exterior nodes of the graph with a question of whether there is a node whose exterior degree doubles. This would not have been possible without this data structure which used the BFS layers to distinguish the distance from the minimum out-degree node as a critical missing link in the SSNC. We will also have to deal with back arcs, but as we will discuss in Section 7 (Back Arcs), that will not be a problem at that point because by then, we will be able to treat them more like exterior neighbors than interior neighbors. With the concepts of interior degree doubling and generalization of load balancing, we will be able to move on to back arcs.
It will seem like back arcs are being avoided in oriented graphs, like traditional transitive triangles were by previous authors [2], [9]. What we will see in section 7 (Back Arcs) is that this is not the case. Although they share the transitive triangle property with interior neighbors, they are more similar to exterior neighbors. These arcs are still neighbors to a node in a different rooted neighborhood. What makes back arcs simpler is that once we have shown the decreasing nature of these rooted neighborhoods, and interior degree doubling, the existence of a back arc will imply a degree doubling node. Thus, we are not avoiding back arcs because of their difficulty, but because we have to prove a lot of concepts before we can use it on the back arcs.
This is the type of scenario, particularly the problematic nature of back arcs and their connection to transitive triangles, that highlighted the need for a more robust and organized framework. The Graph Level Order directly addresses this by first partitioning the graph into a sequence of rooted neighborhoods about the minimum out-degree node $v_{0}$. Within this organized framework, we are able to differentiate between different types of transitive triangles, those involving interior arcs, and those involving back arcs. The critical properties that are enforced by the Graph Level Order (e.g., by ensuring that rooted neighborhoods decrease in size, will be shown in Section 6), will allow us to transform these potentially problematic arcs into instances where the node $w$ has a second neighborhood larger than its first neighborhood. This will make $w$ a degree doubling node, consistent with the conjecture.
This deep understanding, revealed through the Graph Level Order, transforms what were previously intractable combinatorial hurdles—such as those posed by transitive triangles—into manageable components, thereby forging a clear path to a complete proof.
5.2 Proofs Using Graph Level Order
We are trying to prove the SSNC by contradiction. That is, we are supposing it is not true and investigating the properties that must be true in a graph. We are now ready to look again at the set-theoretic aspects of the SSNC. First, though, we need to understand what the set of interior and exterior elements represents.
Lemma 5.2.
Interior Cover Lemma Let $G$ be an oriented graph and $u\in G$. Suppose that for every $x\in N^{+}(u)$, there exists a $v\in N^{+}(u)$ such that $u\to v\in G$ and $v\to x\in G$. Then:
$N^{+}(u)=\bigcup_{v\in N^{+}(u)}\text{int}(u,v)$
Proof.
Recall that
$\text{int}(u,v)=N^{+}(u)\cap N^{+}(v)$
for any $v\in N^{+}(u)$, where $u$ is a parent of $v$.
First, let $x\in N^{+}(u)$. By the premise, there exists $v\in N^{+}(u)$ such that $u\to v\in G$ and $v\to x\in G$. Thus, $x\in N^{+}(v)$, and $x\in\text{int}(u,v)$. Hence,
$N^{+}(u)\subseteq\bigcup_{v\in N^{+}(u)}\text{int}(u,v).$
Conversely, if
$x\in\bigcup_{v\in N^{+}(u)}\text{int}(u,v),$
then there exists $v\in N^{+}(u)$ such that
$x\in\text{int}(u,v).$
Therefore,
$x\in N^{+}(u),$
and
$\bigcup_{v\in N^{+}(u)}\text{int}(u,v)\subseteq N^{+}(u).$
Combining both inclusions, we conclude $N^{+}(u)=\bigcup_{v\in N^{+}(u)}\text{int}(u,v)$. ∎
Lemma 5.2 (Interior Cover Lemma) demonstrates that if every child out-neighbor $x$ of a parent $u$ is the tail of a directed triangle with $u$ as the source, then the out-neighborhood of $u$ can be expressed as the union of the intersections of the out-neighborhoods of $u$ and its out-neighbors. This implies that the out-neighbors of $u$ share the workload of covering the out-neighborhood. Instead of $u$ needing to reach each of its out-neighbors directly, the existence of these transitive triangles ensures that each out-neighbor $x$ is also reached by another out-neighbor $v$. This effectively distributes the responsibility of reaching out-neighbors among the nodes in $N^{+}(u)$, preventing any single node from bearing the entire burden. This distribution can help prevent the node $u$ from becoming a degree-doubling node, provided there are sufficient nodes in $N^{+}(u)$ to handle the distributed responsibility.
Lemma 5.3.
Exterior Cover Lemma Let $G$ be a graph, and let $u\in G$. Suppose that for every vertex $x$ at distance two from $u$, there exists a $v\in N^{+}(u)$ such that $v\to x\in G$. Then:
$N^{++}(u)=\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)$
Proof.
Recall that for any $v\in N^{+}(u)$, the exterior set is defined as
$\text{ext}(u,v)=N^{++}(u)\cap N^{+}(v)$
Intuitively, the set $\text{ext}(u,v)$ represents the portion of $u$’s second out-neighborhood that is reachable on a path through a specific out-neighbor $v$.
Let $x\in N^{++}(u)/N^{+}(u)$. By assumption, there exists some $v\in N^{+}(u)$ such that $x\in N^{+}(v)$. Then:
$x\in N^{++}(u)\cap N^{+}(v)=\text{ext}(u,v)$
So
$x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v).$
Thus,
$N^{++}(u)\subseteq\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)$
Conversely, suppose
$x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v).$
Then there exists $v\in N^{+}(u)$ such that $x\in\text{ext}(u,v)$, which means:
$x\in N^{++}(u)\cap N^{+}(v)$
Hence, $x\in N^{++}(u)$.
So:
$\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)\subseteq N^{++}(u)$
Combining both directions, we conclude:
$N^{++}(u)=\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)$
Lemma 5.3 (Exterior Cover Lemma) demonstrates that if every second out-neighbor $x$ of $u$ is the tail of an arc from an out-neighbor $v$ of $u$, then the second out-neighborhood of $u$ can be expressed as the union of the exterior out-neighborhoods. This condition implies that the out-neighbors of $u$ collectively cover the second out-neighborhood. In other words, every node in $N^{++}(u)$ is reachable from at least one node in $N^{+}(u)$ via a single arc. This ensures that the responsibility of reaching the second out-neighbors is distributed among the first out-neighbors, preventing $u$ from bearing the entire burden alone.
What’s happening is that we have reformed the conjecture without even introducing the Decreasing Neighborhood Sequence Property (DNSP). Rather than talking about second out-neighbors being greater than first, these Lemmas 5.2 and 5.3 are saying that in order for the SSNC to fail at a node, we need to have load balancing. This means, first, that we cannot have an out-neighbor with 0 out-degree in the first out-neighborhood, i.e., we must have cycles present in that rooted neighborhood.
Next, we will prove that any oriented graph can be represented in a Graph Level Order. We will need to demonstrate how the graph can be systematically partitioned into rooted neighborhoods while preserving the Graph Level Order properties. This includes maintaining a total order of the rooted neighborhood, allowing for the comparison of any two nodes, establishing a total order of those nodes within the rooted neighborhood, and establishing interior and exterior out-neighbors.
Theorem 5.1.
Graph Level Order Representation of Oriented Graphs without Back Arcs Given an oriented graph $G=(V,E)$ without back arcs, $G$ can be represented by a Graph Level Order.
Proof.
We aim to show that $G$ can be systematically partitioned into rooted neighborhoods that satisfy the key properties required by the Graph Level Order structure:
We select a node $v_{0}\in G$ of minimum out-degree. This node will serve as the root for defining rooted neighborhoods based on graph distance.
Define rooted neighborhoods $R_{i}=\{u\in G\mid\text{dist}(v_{0},u)=i\}$ for $i=0,1,2,\ldots,k$. Here, $\text{dist}(v_{0},u)$ is the length of the shortest directed path from $v_{0}$ to $u$. Since $v_{0}\in R_{0}$, each $R_{i}$ collects all nodes at distance exactly $i$ from $v_{0}$.
By assumption, $G$ contains no back arcs (arcs going from a node in $R_{i}$ to a node in $R_{j}$ with $j<i$). Therefore, the neighborhoods ,$R_{1},\ldots,R_{k}$ form a natural total order:
$R_{0}<R_{1}<\ldots<R_{k}$
where $u<v$ if and only if $\text{dist}(v_{0},u)<\text{dist}(v_{0},v)$. This total order arises because the distance function respects strict inequality between layers, and no arcs ”go backward” in terms of distance.
Consider any two distinct nodes $u,v\in R_{i}$. Since they share the same distance from $v_{0}$, distance alone cannot order them. To establish a total order within each $R_{i}$, introduce a secondary ordering metric such as:
  • Lexicographic order of adjacency lists,
  • Comparing out-degree values,
  • Or any consistent tie-breaking scheme on node identifiers.
This ensures any two nodes within the same rooted neighborhood can be compared, completing the total order on the entire vertex set $V$.
Because the graph is oriented and contains no back arcs, arcs only go from nodes in $R_{i}$ to nodes in $R_{i+1}$ or within the same $R_{i}$.
Thus, the ordering respects the graph structure, and comparisons are consistent across rooted neighborhoods.
For $u\in R_{i}$ Interior neighbors are defined for $v\in N^{+}(u)\subseteq R_{i+1}$
$\text{int}(u,v)=N^{+}(u)\cap N^{+}(v)$
Exterior neighbors are defined for $v\in N^{+}(u)$: Nodes reachable from $u$ in two steps but not directly from $u$, and which are reachable from $v$:
$\text{ext}(u,v)=N^{++}(u)\cap N^{+}(v)$
Consider an arc $u\to v\in G$, with $v\in R_{i+1}$. This makes $u$ a parent of $v$. Now consider the arc $v\to w\in G$, where $w\in N^{+}(v)$. There are three possible cases. Either $u\to w\in G$ or $u\to w\notin G,\text{ where }w\in R_{i}$ or $u\to w\notin G\text{ where }w\in R_{i+1}$.
Case 1: If $u\to w\in G$, then $w\in\text{int}(u,v)$.
Case 2: If $u\to w\notin G$, but $w\in R_{i+1}$, then $w\in\text{ext}(u,v)$.
Case 3: If $u\to w\notin G$ and $w\in R_{i+2}$, then $w\in\text{ext}(u,v)$.
The partition ${R_{0},R_{1},\ldots,R_{k}}$ forms a leveled hierarchy of rooted neighborhoods, with a total order on rooted neighborhoods and on nodes within each neighborhood. Interior and exterior neighborhoods are well-defined, preserving the Graph Level Order properties. Hence, under the assumption of no back arcs, $G$ can be represented by a Graph Level Order.
This completes the proof. ∎
What the Graph Level Order does is transform the oriented graph into an ordering of rooted neighborhoods with inner-metrics. This ordering is more than just representative. The rooted neighborhoods help to organize the nodes by giving them location identifiers. What this suggests is that each rooted neighborhood is totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in the 2-dimensional plane. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$. Now, each rooted neighborhood has a proper place, with its nodes inside it. These rooted neighborhoods are similar to hyper-nodes in that context since they were defined as induced subgraphs based on the shortest path. In this context, the exterior edges act as hyper-edges between the rooted neighborhoods.
With nodes now given these location identifiers, we can now think of the SSNC in terms of interior neighbors, exterior neighbors and back arcs instead of simply first and second neighbors. This is a more convenient thing because the new terms allow for a partition of a node’s degree (Lemma 5.1 (Partition of Node’s Degree)), whereas the older terms do not. Moreover, there have been problems with concepts like transitive triangles and double counting in the SSNC. We will show in the Lemma 5.2 (Classification of Transitive Triangles in the Graph Level Order) that by giving these nodes identifiers, we can distinguish different types of transitive triangles. This allows us to understand different ways to treat ’problematic’ transitive triangles (i.e. ones that have back arcs), vs ones that help us construct our data structure.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each rooted neighborhood maintains its place in the total ordering. It is a bold assumption to claim that oriented graphs have no back arcs, though. Oriented graphs come in all shapes and sizes. As we will cover in section 7 (Back Arcs), the disjoint nature of the rooted neighborhoods tells us that if any two neighborhoods are ever not disjoint, this is both a necessary and sufficient condition for the existence of a back arc.
Section 7 (Back Arcs) will look into back arcs in more depth. We will prove there that when these arcs do exist, we are not able to construct a Graph Level Order. Instead, we are able to immediately find a degree-doubling node. That is, a back arc immediately leads us to a node whose second out-neighborhood is at least as large as its first out-neighborhood. This means that back arcs do not hurt with the development of our algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a Graph Level Order, as we will have already found the degree-doubling node(s).
Next, we will follow up with our talk about transitive triangles. Now that we know that every oriented graph can be represented in a Graph Level Order, Theorem 5.2(Classification of Transitive Triangles in the Graph Level Order) gives every possible type of transitive triangle that can be represented in a Graph Level Order.
Theorem 5.2.
Classification of Transitive Triangles in the Graph Level Order In any oriented graph $G$ represented via a Graph Level Order rooted at a minimum out-degree node, every transitive triangle falls into exactly one of six structural types based on the position of nodes in rooted neighborhoods and the direction of arcs.
Proof.
We classify all transitive triangles by examining the configuration of the first two arcs and then determine whether a valid third arc can close the triangle transitively.
Case # Exterior # Back # Interior Triangle Possible?
1 0 0 3 Yes
2 0 1 2 No
3 0 2 1 Yes
4 1 0 2 Yes
5 1 1 1 Yes
6 1 2 0 Yes
7 2 0 1 Yes
8 2 1 0 No
Table 1: Enumeration of possible combinations of arc types forming transitive triangles.
We now describe the realizable cases:
  • Case 1 (Interior Triangle): All three nodes lie in the same rooted neighborhood $R_{i}$. Arcs: $x\to y,x\to z,y\to z\in G$. All arcs are interior.
  • Case 2 (Invalid): One back arc and two interior arcs. If both $y$ and $z$ lie in $R_{i}$ and one arc points backward, transitivity cannot hold unless $x$ is in $R_{i}$, which contradicts the rooted structure. Hence, this is not possible.
  • Case 3 (Back Arc Triangle I): Two back arcs and one interior arc. Here, both $x,y\in R_{i}$, and $z\in R_{i+1}$. All arcs are: $x\to y,z\to x,z\to y$. This forms a valid triangle.
  • Case 4 (Interior-Exterior Triangle): One interior arc and two exterior arcs. A common parent $x\in R_{i}$ maps to children $y,z\in R_{i+1}$ with $x\to y,x\to z,y\to z\in G$. Standard configuration for load balancing, defining the interior nodes.
[…]
  • Case 6 (Back Arc Triangle III): One exterior arc and two back arcs connecting three different rooted neighborhoods $R_{i}$, $R_{i+1}$ and $R_{i+2}$, e.g., $x\to y,z\to x,z\to y$, where $x\in R_{i}$ $y\in R_{i+1}$ and $z\in R_{i+2}$ forms a valid triangle.
  • Case 7 (Exterior Arcs Case 2:): Two exterior arcs and one interior arc, e.g., $x\to z,y\to z,x\to y$, where $x,y\in R_{i}$, $z\in R_{i+1}$.
  • Case 8 (Invalid): One back arc and two exterior arcs. The assumption of two exterior arcs implies $x\to y,x\to z)$, where $x\in R_{i}$ and $y,z\in R_{i+1}$. This third arc cannot be $y\to x$ or $z\to y$ since $G$ is an oriented graph. The other option that would make this a transitive triangle would be the arcs $y\to z$ or $z\to y$, neither of which are back arcs since $y$ and $z$ are in the same rooted neighborhood.
Interior Triangle Back Arc Triangle (Case I) Interior-Exterior Triangle
[…]
Back Arc Triangle (Case II) Back Arc Triangle (Case III) Exterior Triangle
[…]
Table 2: Examples of transitive triangle types observed in the Graph Level Order structure: Nodes are given identifiers based on their neighborhoods. From there, arcs are determined as interior, exterior, and back arc depending on the two node endpoints. We see six possible transitive triangle cases. Unlike exterior and interior arcs, back arcs are drawn with dotted lines.
Since all structurally valid triangles fall into one of these categories, and we have excluded all others through case exhaustion, the theorem follows. ∎
This further illustrates the power of the Graph Level Order. More importantly, we see transitive triangles arise naturally in domains such as social media, to represent friendship recommendations which is not a Euclidean, These social media platforms are often structured around principles like triadic closure—the tendency for two individuals with a mutual friend to form a connection—leading to the frequent emergence of transitive triangles.
What Theorem 5.2 demonstrates is that transitive triangles present a non-uniform challenge to the (SSNC). There is no one-size-fits-all approach to handling them:
  • Back arc triangles must be treated using one class of structural arguments, exploiting the fact that they are going to previous, larger neighborhoods.
  • Interior triangles require a second method, often leveraging cycles, load balance and degree-doubling within a single neighborhood level.
  • Exterior triangles necessitate a third strategy, which would focus on traversing through the remainder of the graph.
This nuanced categorization and the ability to reason about these cases independently and precisely would not be possible without the Graph Level Order, which partitions the graph by both distance and degree in a way that preserves local structure and directional influence. The Graph Level Order thus enables the decomposition of complex global conjectures into manageable local conditions, each of which can be analyzed with tools tailored to its structural type.
The presence or absence of these triangles can reveal critical information about the graph. Crucially, by categorizing these transitive triangles into these distinct types, we are able to design targeted strategies for analyzing and responding to each case. This dissection enables the development of algorithms that move beyond treating transitive triangles as a homogeneous phenomenon. Instead, we can apply a fine-grained, type-specific reasoning to each instance.
Next, we will show that one of the other benefits of this data structure is mathematical induction via well ordering. Mathematical induction is a proof technique that uses inference to prove statements. Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element. This well-ordering property allows us to use mathematical induction on the rooted neighborhoods, where the base case is the minimum out-degree node $v_{0}$, and the inductive step involves reasoning about the properties of $R_{i+1}$ based on the properties of $R_{i}$.
Lemma 5.4.
Graph Level Order Supports Induction In any oriented graph, without back arcs, the minimum out-degree node $v_{0}$ within a Graph Level Order structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
First, remember that we are under the assumption that we do not have back arcs. We have already shown that under this assumption, the rooted neighborhoods in a Graph Level Order form a total order. Further, they are a discrete total order. By definition, a set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order, where each element has a distinct next element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a smallest element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. Thus, the total order defined by the distance from $v_{0}$ on the set of rooted neighborhoods $R_{1},R_{2},\ldots$ constitutes a well ordering. ∎
This becomes an important concept because now, in addition to the nodes, we also have the rooted neighborhoods as subsets of $G$. These rooted neighborhoods are totally ordered and thus can also be used as a means of mathematical induction. This gives us several options of how to proceed with our proofs, but the foundation will be the Graph Level Order, which orders the nodes based on their distance from $v_{0}$-the minimum out-degree node.
Not all oriented graphs possess the properties of being well-ordered. In fact, we have been working under the assumption that there are no back arcs precisely because they are one such exception. These back arcs prevent a well ordering, which are what we need for mathematical induction. However, under the assumption of no back arcs, we do arrive at a well ordering. The proof of Lemma 5.4 established that every element has a uniquely defined next element. Back arcs would obscure this
However, within the specific context of the SSNC and the Graph Level Order we have developed here, we posit that any oriented graphs that do not fit into this well-ordering immediately give rise to a degree-doubling node. This will be discussed more in Section 6 (The Decreasing Neighborhood Sequence Property) and Section 7 (Back Arcs) and how those two interact.
What we will see is that the DNSP initially acts on the graph at the individual node level. This establishes the condition that if a node does not satisfy the conjecture (i.e., $|N^{++}(u)|<|N^{+}(u)|)$, and the second out-neighborhood is smaller than the first. We will then show that the DNSP rises to the rooted neighborhood level, where at rooted neighborhood $i$, it cannot have more nodes than rooted neighborhood $i+1$ for $i\geq 1$. What back arcs do in this context is create a link from a smaller rooted neighborhood to a larger rooted neighborhood (i.e. a rooted neighborhood $i$ to a rooted neighborhood $j<i$, forcing the nodes in the larger rooted neighborhood $i$ to violate the DNSP constraints when considering their connections to the smaller, earlier rooted neighborhood $j$. The core proof idea that is necessary for the back arcs to hold is that of interior degrees doubling. This allows for a traversal algorithm of exterior degrees to take place. This will be presented in Section 6 (The Decreasing Neighborhood Sequence Property),
While the SSNC might seem like a degree vs. distance problem, the Graph Level Order doesn’t pick a side. Instead, it uses degree to establish a foundation, via the minimum out-degree node. We are then able to use that node to partition the graph, and take advantage of the ordered structure created by distance. This ordered distance is key to the Graph Level Order’s analytical power, but it wouldn’t exist in this form without the initial degree-based choice of the anchor. Therefore, the Graph Level Order achieves a balance by making them interdependent rather than mutually exclusive.
While the Graph Level Order data structure powerfully fine-tunes our intuition and aids in the discovery process, its utility must be carefully balanced with the demands of mathematical rigor. The photographs and illustrations offer an understanding of the general behavior of the nodes, but we have to be careful with that. Our mathematical proofs must not rely on this visual intuition alone. Each step must proceed from a logical step, grounded in definitions, axioms, theorems and proven results. This requires us to sometimes omit one of the very concepts that makes this Graph Level Order special, its total order and how this is so easy to be represented pictorially. This is a necessary journey as the Graph Level Order begins a transition from a simple data structure concept, to a powerful tool capable of providing the means to prove the SSNC.
6 The Decreasing Neighborhood Sequence Property
6.1 Introduction
Having laid the groundwork for the Graph Level Order, we are ready to begin using it. The next and most important concept that we will define that will guide our research is the Decreasing Neighborhood Sequence Property (DNSP). With that and the Graph Level Order, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a Graph Level Order, we can assume we are working in this environment.
The DNSP explains the nature of how the sizes of the rooted neighborhoods decrease as we get further from $v_{0}$. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it, though. As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase since these degrees must go somewhere. This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum out-degree node, $v_{0}$, serves as a catalyst and driving force of this process. At the other end of the spectrum, the Decreasing Neighborhood Sequence Property is the bottleneck limiting the expansion of the graph’s out-degrees.
The case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) is important because it provides information about what a counterexample of the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, when the arcs are oriented in the same direction, a cycle of order three is created among the rooted neighborhood adjacent to $v_{0}$. If we look at this situation further, we see in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will consider the assumption of other nodes in the graph to not have their degrees double.
However, we will need to define this new concept formally to proceed with this line of reasoning. This property is attributed to a node in $G$ whose out-degree does not double. There are instances where $v_{0}$ lacks its out-degree double, as we have seen.
Definition 6.1.
For a node $u\in G$ in an oriented graph $G$, we say that $u$ has the Decreasing Neighborhood Sequence Property (DNSP) if the size of its second out-neighbors is strictly smaller than the size of its first out-neighbors, i.e., $|N^{++}(u)|<|N^{+}(u)|$.
Figure 9: This figure illustrates the Decreasing Neighborhood Sequence Property, showing a node 0 with five first out-neighbors and four second out-neighbors Thus, the size of rooted neighborhood 1 is greater than the size of rooted neighborhood 2.
There is a fundamental result in graph theory stating that every Directed Acyclic Graph (DAG) must contain a sync [19], or a vertex with no outgoing edges. This insight directly inspired the development of the Decreasing Neighborhood Sequence Property. Here, we formulate a hypergraph where each of the rooted neighborhoods $R_{0},R_{1},\ldots,R_{k}$ can now be re-interpreted as a hyper-node. The exterior arcs between rooted neighborhoods can then be viewed as hyper-arcs. This yields a hypergraph structure. When there are no back arcs, this hypergraph becomes a Hyper Directed Acyclic Graph (Hyper-DAG). As such, the rooted neighborhoods admit a total ordering from the minimum out-degree node $v_{0}$ through a terminal node that acts as a sink in the hypergraph. This provides both a theoretical foundation and a practical framework for understanding degree growth and influence propagation in oriented graphs.
The proof by contradiction implied that there must be some inherent organization at the level of rooted neighborhoods. From standard graph theory, we knew that any directed acyclic graph (DAG) must contain a sync. By the data structure construction, all nodes are pointing forward toward that sync and away from the root node through exterior arcs. While each individual node satisfies the Decreasing Neighborhood Sequence Property (DNSP), our interest lies in whether this local condition produces a global effect on the hypergraph level. To avoid having a node’s degree double, there must be fewer exterior neighbors forward, and thus fewer exterior arcs and hyper arcs forward to the following neighborhoods. This scarcity is what ultimately causes the sizes of neighborhoods to decrease.
The SSNC asks for a node whose second out-neighborhood is at least as large as its first. Our strategy is to prove the conjecture by contradiction. We assume the Decreasing Neighborhood Sequence Property holds for all nodes in $G$. The definition of the DNSP is a direct negation of what the conjecture asks for. This assumption allows us to shift our focus from directly searching for a degree-doubling node to analyzing the graph’s global composition. One notable change is that the DNSP calls for the second out-neighborhood to be strictly less than the first. What we would like to do is set up a decreasing exterior condition, which could prove strong. We anticipate that this exploration will reveal constraints on the graph’s arrangement that will ultimately lead to a contradiction. The problem with this approach is that even with visually appealing results, they necessitate mathematical rigor.
Two competing new ideas have been presented: the Decreasing Neighborhood Sequence Property (DNSP) and the minimum out-degree node. A crucial starting point for our approach and out-degree values is the minimum out-degree node $v_{0}$. The minimum out-degree node acts as the anchor by being chosen as the starting point. By selecting it as a minimum out-degree node, every other node in the graph has one other constraint that must be met based on $v_{0}$. Similar to Constraint Satisfaction Programming (CSP), as we move further from $v_{0}$, more requirements are being placed on these nodes in order to keep the additional nodes from becoming degree-doubling nodes. Each rooted neighborhood $R_{i}$ gives us more parent-child relationships, and those new children have the load balancing responsibilities of keeping their parents and everyone else in the hierarchy from being overwhelmed by the metaphorical load. The node $v_{0}$ is acting as a driving force, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DNSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network [10], a role precisely filled by the DNSP. Remember, we are attempting to prove this conjecture by contradiction. That is, we are assuming that no node in the graph is a degree-doubling node. While the node $v_{0}$ places minimum out-degree constraints on each node, this bottleneck constraint limits how much of that can go towards the exterior degrees of these nodes, and thus how many exterior neighbors these nodes can have. Thus, the minimum out-degree node $v_{0}$ and the DNSP property together function as a catalyst and bottleneck, respectively. We will see how they are in competition with one another throughout the rest of this paper.
The DNSP formalizes what it means to investigate nodes on a local level. It is, in essence, the node-level negation of the SSNC: a node has DNSP if its second out-neighborhood is strictly smaller than its first. Our focus is to identify what elements of a graph cause a node to have this property. We would also like to determine how many such conditions can exist before that in itself causes the load to collapse. Through this lens, we can better understand these local obstacles to the conjecture. We would like to see how far we can extend this. The following lemma tells us the methodology behind this approach. All we need is one node to fail the DNSP conditions in order for it to become a degree doubling node.
Lemma 6.1.
Decreasing Neighborhood Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the Decreasing Neighborhood Sequence Property. Then this node will have its out-degree at least double in $G^{2}$.
Intuitively, what this lemma is saying is that nodes that violate the DNSP will automatically satisfy the conjecture and no longer need to be investigated. Our proof structures will then focus on the more challenging case where every node in $G$ has the DNSP. This gives us a strategy for our proofs. We will set them up as proof by contradiction. Once we have shown a node has violated the DNSP, then this lemma is implied and we have found a degree doubling node.
In this work, we assume that the graph is oriented, unweighted, and without back arcs. We also assume that all nodes in this graph have the Decreasing Neighborhood Sequence Property. What that means is that every node’s second out-neighborhood is strictly smaller than its first out-neighborhood. This is still the setup for a proof by contradiction. In the previous section, we showed that any unweighted oriented graph without back arcs could be represented in a Graph Level Order. Instead of allowing nodes to be placed anywhere across the plane, we will take advantage of this structure here. The nodes will be partitioned into rooted neighborhoods with interior arcs to the current rooted neighborhood and exterior arcs to the next rooted neighborhood. For the remainder of this paper, when we speak of oriented graphs, the underlying assumption is that it is an oriented graph in a Graph Level Order.
Lemma 6.2.
Interior Load Balancing Lemma Suppose we have an oriented graph $G$, with minimum out-degree node $v_{0}$. Then, every node $x$ in the neighboring rooted neighborhood $R_{1}$ must have an interior degree of at least 1. That is, for $x\in R_{1}$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the Decreasing Neighborhood Sequence Property if
$|N^{+}(x)|>|N^{++}(x)|.$
By the definition of an interior arc,
$\text{int}(v_{0},x)=N^{+}(v_{0})\cap N^{+}(x)$
A node $x\in R_{1}$, will have exterior out-neighbors in $R_{2}$, making them second out-neighbors of $v_{0}$. If $x$ has zero interior arcs, then $x$ must have $\delta=d^{+}(v_{0})$ exterior arcs. This means that the size of $v_{0}$’s second neighborhood, $|N^{++}(v_{0})|$ would be at least $|N^{+}(v_{0})|$, violating the DNSP condition for $v_{0}$. This implies that $x$ must have at least one interior arc to another node in $R_{1}$. By definition, $v_{0}$ has $d^{+}(v_{0})$ first out-neighbors. This means that
$N^{++}(v_{0})\geq N^{+}(v_{0}).$
To avoid this doubling of $v_{0}$, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
$|\text{int}(v_{0},x)|\geq 1$
This is true for the out-neighbors of every node in $R_{1}$. This ensures that every node in $R_{1}$ has an interior degree of at least one. Hence, every node $x\in R_{1}$ satisfies
$|\text{int}(v_{0},x)|\geq 1.$
Thus, for all $x\in R_{1}$,
$|\text{int}(v_{0},x)|\geq 1.$
Lemma 6.2 (Interior Load Balancing Lemma) is important because we can see some properties in generally oriented graphs that may coincide with degree-doubling. We see that when all of the out-neighbors of the minimum out-degree node have degree 1, then it will prevent that minimum out-degree node from having its degree doubled. This is a generalization of Lemma 4.3 (Minimum Out-Degree 3 with Out-Neighbors 1).
A fundamental concept in directed graphs is that when all nodes have out-degree 1, there must be a cycle. In the context of the Graph Level Order, this simple case provides an intuitive understanding of our load balancing mechanism. To prevent the minimum out-degree node, $v_{0}$, from becoming a degree-doubling node, the first rooted neighborhood, $R_{1}$, must exhibit internal connections with every node in $R_{1}$ participating in a cycle. This cyclic structure allows the nodes in $R_{1}$ to share the load, reducing the number of distinct out-neighbors $v_{0}$ needs at distance 2. This principle of local interconnectedness, where nodes within a rooted neighborhood form cycles to manage the flow of connections, is central to the Graph Level Order’s ability to prevent degree doubling at each level. In particular, the presence of cycles within $R_{1}$ allows nodes to reuse neighbors as second out-neighbors of $v_{0}$, reducing the total number of distinct second out-neighbors and preventing degree doubling.
Lemma 6.3.
Exterior Load Balancing Lemma Suppose we have an oriented graph $G$. If $x$ in rooted neighborhood $R_{i}$ has the Decreasing Neighborhood Sequence Property, then for all $y\in N^{+}(x),|\text{ext}(x,y)|<d^{+}(x)$. That is, for any out-neighbor $y$ of $x$, the number of exterior out-neighbors of $x$ which are reached through $y$ must be less than the out-degree of $x$.
Proof.
Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$. By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$. Thus,
$|\text{ext}(x,y)|\subseteq N^{++}(x).$
Then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x)$
This means that the out-degree of $x$ at least doubles on the graph $G^{2}$. This contradicts the assumption that $x$ has the DNSP. This must be true for all $x$ in $G$, so we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
Lemma 6.3 (Exterior Load Balancing Lemma) is very similar to Lemma 6.2 (Interior Load Balancing Lemma), with the notable exception that Lemma 6.3 refers to any node with the DNSP, whereas we only made Lemma 6.2 true for the minimum out-degree node. This generalization emphasizes the pervasive role of exterior neighborhood structure in maintaining the DNSP throughout the graph. Lemma 6.3 may seem like a stronger lemma because it applies to more nodes, but because a node’s degree is partitioned by interior and exterior neighbors, having an exterior neighbor less than the node’s overall degree is not as strong as it seems. The stronger lemma is what will come later, Lemma 6.6 (Decreasing Exteriors), which states that the size of these exteriors decrease as we move about a path in the Graph Level Order.
The results presented in 6.2 (Interior Load Balancing Lemma) and 6.3 (Exterior Load Balancing Lemma) are closely related to Daamouch’s [6] Lemma 1.1 and 1.2. There, he demonstrated that a minimum out-degree vertex with no cycles in its induced graph has a degree that effectively doubles. He also showed by the same reasoning that the minimum degree $\delta$ of an oriented graph must be at least 3, that the first neighborhood the minimum out-degree node will be the disjoint union of directed triangles and that the cardinality of the second neighborhood is $\delta-1$.
This similarity to Daamouch’s work shows the power in transitive triangles. As mentioned, other researchers have connected transitive triangles to the SSNC. We have already mentioned the work of Brantner et al. in first bringing them to the conjecture. This paper by Daamouch did extensive work in studying transitive triangles in the context of $m$-free graphs and anti-transitive oriented graphs. He was able to prove that the conjecture holds for certain classes of these graphs-namely all classes of 5-anti-transitive graphs. He also found that these classes had at least two Seymour vertices.
Despite the overlap, this paper will be moving in a different direction from Daamouch and others with whom we share similarity. We have defined fundamental concepts like the rooted neighborhoods, the Graph Level Order and the Decreasing Neighborhood Sequence Property, which will all help build towards a global search solution. The similarity of the proofs we have presented so far was first because this research has developed over a number of years. Second, these proofs are all building a data structure and an algorithm brick by brick. Later proofs will depend on these results. And while the work of Brantner et al., Daamouch and others is similar, it is not quite the same, as their work did not have terms like exterior neighbor or the DNSP defined. As such, using only references instead of explicitly showing that these proofs could be done within the Graph Level Order would mean potentially missing a case that we could have explained.
For the remainder of this section, we operate under the assumption that every node $u$ in the oriented graph $G$ satisfies the Decreasing Neighborhood Sequence Property: $|N^{+}(u)|>|N^{++}(u)|$.
6.2 Neighborhood Density
The power of the Decreasing Neighborhood Sequence Property (DNSP) and the Graph Level Order goes well beyond the above two lemmas that have been mentioned in literature in some form, as we will prove. In fact, the above lemmas can be generalized, but first we need to show that the DNSP can show more about the overall graph. Under the standing assumption that every node satisfies DNSP, $|N^{++}(u)|<|N^{+}(u)|\forall u\in G$, we proceed to explore Neighborhood Density. This concept will extend beyond pairwise neighborhood comparisons to offer insights into the broader structural constraints imposed by DNSP on the graph.
The earlier lemmas (Interior and Exterior Load Balancing) are specific instances of how DNSP restricts node neighborhoods locally. However, DNSP’s implications reach further, shaping the overall topology and connectivity patterns—especially when analyzed in the context of the rooted neighborhoods and the Graph Level Order.
One additional thing that makes the rooted neighborhoods important is that they are each their own instance of the SSNC. What we mean by that is that we can use a divide and conquer approach, and individually isolate the rooted neighborhood $R_{i}$ from the rest of the graph. The SSNC would apply to the rooted neighborhood $R_{i}$. This rooted neighborhood $R_{i}$ is also an oriented graph and so it can be put into a Graph Level Order. This gives a relationship between the nodes in the rooted neighborhoods and different ways they can interact with nodes outside the rooted neighborhoods. Because these nodes have been built within this data structure, we do not need to understand them in terms of any oriented graph. Instead, we have more of a question of the preventative measures we have been taking to avoid degree doubling nodes.
We have discussed how nodes are required to do load balancing to support parent nodes. What nodes also need to do is have their ’interior’ degrees double. What we mean by this is that, in the absence of back arcs, we know that a node’s degree equation can be split into interior and exterior degrees. In the rooted neighborhood $R_{i}$, the node $u_{i}$ will then have the out-degree of $\text{int}(u_{i-1},u_{i})$, where $u_{i-1}\in R_{i-1}$ is a parent of $u_{i}$. The statement that the interior degree doubles is just saying that the SSNC holds for the node $u_{i}$ when looking at the interior degree as its out-degree in $R_{i}$.
This gives us a way backward and a way forward. The path backward is the nodes in a current neighborhood $R_{i}$ using load balancing to help their parents to avoid becoming degree doubling nodes. The way forward is through interior degree doubling. With this, a node’s degree equation, which was reduced to interior degree and exterior degree, is simply focused on exterior degree going forward. Once we have proven that the interior degree doubles, we will be able to simply search for nodes whose exterior degree doubles as well. This will open the door for our path-finding traversal through exterior nodes.
Before we discuss the interior degree doubling lemma, let’s introduce it with an example. We have seen with Lemma 6.2 (Interior Load Balancing) that we need to have a cycle in the first rooted neighborhood, but this does not tell us anything about the remaining neighborhoods. The next Lemma, 6.3 (Exterior Load Balancing)lets us know that for every other node the exterior degree must be less than the degree of that node. This gives us a relationship going forward about the decreasing nature of the rooted neighborhoods.
Example 6.1.
Figure 10: An illustration of the minimum degree 7 case with three rooted neighborhoods, all surrounding a minimum out-degree node. Interior arcs are within the neighborhoods and exterior arcs are between the neighborhoods.
Our assumption in this example is that all nodes have the DNSP. We begin with the node $v_{0}$. Because of the DNSP and Lemma 6.2 (Interior Load Balancing), know that every node $u\in R_{1}$ will have $\text{int}(v_{0},u)\geq 1$. Similar to Example 4.2, the cycle(s) in the rooted neighborhood $R_{1}$ allows the nodes in $R_{1}$ to have less than 7 exterior neighbors, while still having their required number of out-neighbors. These exterior neighbors for the nodes in $R_{1}$ are the second neighbors for $v_{0}$. Since there are less than 6 nodes in $R_{2}$, we see that $v_{0}$ is not a degree doubling node, since $6<7$. The situation where every node in $R_{1}$ connects to more than one interior arc means that the exterior degree of every node in $R_{1}$ simply needs to be less than 6 (Every node in $R_{1}$ having 2 interior arcs and 5 exterior arcs would suffice, every node in $R_{1}$ having 3 interior arcs having 3 interior arcs and 4 exterior arcs would suffice as well. Both these situations lead to the sequence simply shrinking faster. As there are only $\binom{7}{2}=21$ arcs, nodes cannot have an average degree higher than 3. ). No matter the situation with the remaining arcs, the cycle in $R_{1}$ is what allows $v_{0}$ to not be a degree doubling node.
Continuing on to the nodes in $R_{2}$, they also have a minimum overall out-degree of at least 7. We want to show that every node in $R_{2}$ will need to have an interior out-degree of at least 2. To see this, consider their exterior neighbors, which map to the rooted neighborhood $R_{3}$. By the definition of exterior neighbors, the exterior neighbors in $R_{3}$ are second neighbors of parent nodes in $R_{1}$ and first neighbors of sibling nodes in $R_{2}$. If a node in $u_{2}\in R_{2}$ has only one interior neighbor in $R_{2}$, then $u_{2}$ must have at least six exterior neighbors in $R_{3}$.
Remember Lemma 5.1 (Partition of Node’s Degree) that states that a node’s degree can be partitioned into its interior out-degree, exterior out-degree and back arcs. We are under the assumption of no back arcs. Next, we notice that for nodes of low degree, as a parent $u_{1}\in R_{1}$ of $u_{2}$, they have their degrees double trivially [20]. Combine this with the fact that this same parent’s exterior degree will now double by exterior degree traversal, because two neighborhoods are the same size $|R_{2}|=|R_{3}|$. That is $|R_{2}|$, the number of first exterior neighbors of $u_{1}$, and $|R_{3}|$, the number of second exterior neighbors of $u_{1}$ are equal in size. By this equality we have that $u_{1}$’s exterior degree doubles. We already showed that $u_{1}$’s interior degree doubles, and since we are assuming that there are no back arcs we have that $u_{1}$’s overall degree doubles. So a node in $R_{2}$ cannot have an interior out-degree of $<2$. This implies that every node in $R_{2}$ must have interior out-degree of at least 2.
Consider the case where a node $u_{2}\in R_{2}$ has a higher interior out-degree than 2. The most balanced degree distribution of the $\binom{6}{2}=15$ possible arcs is (3, 3, 3, 2, 2, 2). Even in this distribution we see that there exists a node with interior out-degree 2, so $R_{3}$ would need to have 5 nodes. More skewed distributions lead to even more nodes with interior out-degree of 2 or less. Consider the distributions (4, 3, 2, 2, 2, 2) and (5, 2, 2, 2, 2, 2). They both lead to even more nodes of degree two. Thus in order to meet the minimum out-degree requirements, there must be at least 5 nodes in $R_{3}$. Five is a prime number, so we know that every node in $R_{2}$ will map to two disjoint cycles in $R_{3}$ of length five. Like in Lemma 6.2 (Interior Load Balancing), these two cycles in $R_{3}$ help prevent the degree doubling of any node in $R_{2}$. Each node in $R_{3}$ is a second neighbor of a node in $R_{1}$. Let us perform calculations on the nodes in $R_{1}$. First, we have already calculated their interior degrees, and because of their low degree, we know that every node in $R_{1}$ will have their degree in $R_{1}$ double. Next, we calculate, their first and second exterior degrees. The first exterior degree of the node $u_{1}\in R_{1}$ will be $|R_{2}|=6$. The second exterior degree of the node $u_{1}\in R_{1}$ will be $|R_{3}|=5$. Because we have $|R_{2}|>|R_{3}|$, the node $u_{1}$ has more first neighbors than second neighbors and $u_{1}$ is not a degree doubling node. The same is true for every node in $R_{1}$.
Next, we move to the nodes in the rooted neighborhood $R_{3}$. Consider, first, the case that some node $v_{3}\in R_{3}$ has an interior degree less than 3. Similar to the $R_{2}$ case, we know that the node $v_{3}$ has a parent $v_{2}\in R_{2}$ that has interior degree of two. What we were able to show by Lemma 4.1 (Minimum Out-Degree $<$ 3) is that nodes with these small degrees will be degree doubling nodes. This means that $v_{2}$ will have its interior degree double. We need to consider the exterior degree of $v_{2}$. Since $\text{int}(v_{2},v_{3})\leq 2$, we have that $\text{ext}(v_{2},v_{3})\geq 5$. By definition, the exterior neighbors that $v_{2}$ has through $v_{3}$ are the second neighbors of $v_{2}$ that are first neighbors of $v_{3}$. This places $\text{ext}(v_{2},v_{3})$ into the next rooted neighborhood $R_{4}$ since every node in $R_{2}$ is a parent of every node in $R_{3}$ by our construction. This implies that $v_{3}$ will have at least five exterior neighbors. This means that $v_{2}\in R_{2}$ of will have five first exterior neighbors and five exterior second neighbors, causing it’s degree to double, which we do not want. This means that we need every node in $R_{3}$ to have interior out-degree 3.
Unfortunately, what we see is that the rooted neighborhood $R_{3}$ requires an oriented graph of 5 nodes, each of degree 3 inside it to support load balancing. This is not possible to construct. To see this, the complete simple graph on 5 nodes has 10 arcs. However, every node with out-degree 3 requires $5\cdot 3=15$ arcs. Thus, we are requiring more arcs than is possible in an orientation of the complete graph of 5 vertices. This means that some node $u\in R_{3}$ will have an exterior degree at least 5. Because of this, we have a violation of the DNSP. By the cycle constructions as we have previously seen, we will be able to have interior degrees doubling, so this violation of the DNSP leads to a node in $R_{2}$ becoming a degree doubling node.
Example 10 just showed how the minimum degree 7 case looks in a Graph Level Order. We also saw how beneficial it is to our strategy of splitting a node’s total degree into its interior and exterior components. Furthermore, we observe that because the interior degree is part of a cycle within the first neighborhood, every node within that cycle will have its interior degree doubled. As a result of this, if we can prove that the exterior degree will also consistently be such that the total second neighborhood of a node is greater than or equal to its total first neighborhood, then we will have an algorithm for the SSNC.
Our next lemma generalizes this example. We will show that our strategy of first dividing a node’s total degree into its interior degree and exterior degree, and then demonstrating that the interior degree will always double, allows us to gain significant ground on the problem. This approach effectively separates the problem into analyzing individual rooted neighborhoods, which are essentially local instances of the (SSNC). Each rooted neighborhood $R_{i}$ is independent of other rooted neighborhoods $R_{j}$, requiring only the information contained within its own nodes and arcs for internal analysis. This is similar to a divide and conquer approach, but without recursion. Knowing that a node’s interior degree has doubled means we only need to examine its exterior neighbors to determine if that node’s overall second neighborhood is at least as large as its first, which is the core of the conjecture.
Lemma 6.4.
Interior Degree Doubling Let $G$ be an oriented graph, and let $v_{0}$ be a node of minimum out-degree in $G$. If nodes at distance $i$ must have degree at least $i)$, then for every rooted neighborhood $R_{i}$ at distance $i$, there exists at least one node whose second interior out-degree is at least as large as its first interior out-degree:
$|N^{++}_{\text{int}}(v)|\geq|N^{+}_{\text{int}}(v)|.$
Proof.
We will provide a constructive algorithm and use properties of permutation groups to show that the condition holds. That is, we will show that we can always map each node in $R_{i}$ to $i$ disjoint cycles. The proof relies on the fact that the internal structure of $R_{i}$ is already known to be a permutation into disjoint cycles of length $\geq 3$. This inherent structure ensures that within $R_{i}$, each node has exactly one outgoing and one incoming edge.
For a node $v\in R_{i}$, its first interior out-degree, denoted $|N^{+}_{\text{int}}(v)|$, refers to the number of nodes $v\to u$ such that $u\in R_{i}$. Similarly, the second interior out-degree of $v$, denoted $|N^{++}_{\text{int}}(v)|$, refers to the number of nodes $w\in R_{i}$ such that there exists a node $u\in R_{i}$ with edges $v\to u\to w$.
We begin this proof with a statement of a simple algorithm and will show that this algorithm is feasible for the Interior Degree Doubling. The purpose of this algorithm is to give an example as a generalization of a load balancing algorithm, and show this is possible. This is not the only such algorithm. Indeed, the number of derangements $D_{n}$ grows at a complexity of $\frac{n!}{e}$. That fraction includes transpositions, or 2-cycles which are outlawed. Removing the number of 2-cycles from this growth rate does not reduce this by much. We outlaw 2-cycles because they are not allowed in oriented graphs and we would like this permutation to transpose back to our graph theoretical framework. To that nature, we have already proved in Section 4 (Initial Lemmas) that all our counter-examples must have minimum out-degree at least 3. This agrees with the minimum degree requirements for a cycle.
Algorithm 1 MapInteriorDegrees
1:A list of nodes $R_{i}$, integer distance $i$
2:Each node in $R_{i}$ is assigned exactly $i$ interior neighbors
3:Initialize a dictionary $assigned\_neighbors$ mapping each node to an empty set
4:Let $n\leftarrow|R_{i}|$
5:for each node $u\in R_{i}$ do
6:     $assigned\leftarrow 0$
7:     for $step=1$ to $n-1$ do
8:         $v\leftarrow R_{i}[(\text{index}(u)+step)\bmod n]$
9:         if $v\neq u\And v\notin assigned\_neighbors[u]$ then
10:              Add $v$ to $assigned\_neighbors[u]$
11:              $assigned\leftarrow assigned+1$
12:         end if
13:         if $assigned=i$ then
14:              break
15:         end if
16:     end for
17:end forreturn $assigned\_neighbors$
Algorithm 2 Given a rooted neighborhood $R_{i}$, this algorithm assigns each node to participate in exactly $i$ interior cycles, simulating repeated interior degree mappings for each level. This construction ensures every node maps to interior neighbors via multiple cycle layers, supporting the interior degree doubling condition.
Example: Consider the neighborhood $R_{i}=Z_{6}=\{1,2,3,4,5,6\}$. One possibility for the algorithm is to return the following two disjoint permutations:
(1 2 3 4 5 6) and (1 3 5)(2 4 6)
Because each node appears once in each cycle representation, these two permutations correspond to two independent mappings. Continuing, since each node has a first neighbor in each cycle, this yields:
$|N_{int}^{+}(v)|=2,|N_{int}^{++}(v)|=2.$
This makes every node in $R_{i}$ a degree doubling node.
Even with limitations on 2-cycles, such a mapping algorithm will allow for the utilization of combinatorial space in $R_{i}$. We can actually look at this as we begin in the rooted neighborhood $R_{i}$.
  • There are $i$ choices for the starting node.
  • There are $i-1$ choices for the second node (we must exclude the first node).
  • There are $i-2$ choices for the third node (we must exclude the first two nodes.
  • $\ldots$
  • There is one choice for the last node (we must exclude all other nodes).
This results in $i!$ possible candidate paths (modulo orientation constraints). This number of such derangements still grows factorially at roughly $\frac{n!}{e}$. This indicates that we still have an abundance of cycle permutations with the 2-cycle restrictions.
The structure of these rooted neighborhoods forms a permutation on the set $R_{i}$, which decomposes into disjoint cycles. Since oriented graphs prohibit 2-cycles, each of these cycles must have a length $k\geq 3$. A key result from group theory (Lagrange’s Theorem) tells us that the order of any subgroup divides the order of the full symmetric group on $R_{i}$, $S_{|R_{i}|}$. These permutations decompose into disjoint cycles whose lengths divide $|R_{i}|$, and the cycles are structure-preserving under composition. Because Lagrange’s theorem partitions $R_{i}$, into subsets that divide $|R_{i}|$, we know that any remaining permutations will always divide the order of the group and no nodes will ever be left without a cycle.
Thus, each node is mapped to a unique successor in its cycle. Therefore, in every cycle of length $k\geq 3$, each node has out-degree 1 and maps to another node in the same cycle, producing a single out-edge within $R_{i}$. The node it maps to will also have a unique out-neighbor since cycles do not repeat nodes. This yields a second-level interior neighbor.
Consider any node $v$ within one of these cycles of length $k\geq 3$. By definition of a cycle within $R_{i}$, $v$ has exactly one outgoing edge $v\to u$ where $u\in R_{i}$. Therefore, $|N^{+}_{int}(v)|=1$. Furthermore, since $k\geq 3$, $u$ must also have exactly one outgoing edge $u\to w$ where $w\in R_{i}$ and $w\neq v$. Thus, $w$ is a unique second-level interior neighbor, implying $|N^{++}_{int}(v)|=1$. Since for every node $v$ in any such rooted neighborhood $R_{i}$, we have
$|N^{+}_{\text{int}}(v)|=1$
and
$|N^{++}_{\text{int}}(v)|=1,$
the condition
$|N^{++}_{\text{int}}(v)|\geq|N^{+}_{\text{int}}(v)|$
is satisfied.
Since such cycle decompositions always exist under the algorithm and orientation constraints, there must be at least one node in every $R_{i}$ that satisfies the interior degree doubling condition.
This result is also aligned with Lemma 4.1 (Minimum Out-Degree $<3$) and related results from [20], confirming that low out-degree nodes within cycles tend to satisfy the degree doubling property. ∎
The positive answer to interior degree-doubling is helpful to the construction of our algorithm. In the absence of back arcs, we can split a node’s degree into interior and exterior degrees. Knowing then that when a node is forced to take upon $i$ cycles in the rooted neighborhood $R_{i}$, that node’s interior degree will have to double means that if that node’s exterior degree also doubles, then that node’s overall degree will double. This gives light to our overall strategy. We will first prove a generalization of the interior load balancing lemma. To do this, though, we need to first prove that the exterior neighbors decrease as we traverse a path. We will assume that all nodes are bound by the Decreasing Neighborhood Sequence Property (DNSP). Then if any node has both its interior and exterior degree double, its overall degree will double, which will lead to a contradiction of the DNSP.
Lemma 6.4 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the rooted neighborhoods as an oriented graph and searched for a node whose degree doubled inside that rooted neighborhood. The proof uses abstract algebra: we can always construct the necessary cycles in the rooted neighborhood $R_{i}$ to make a node’s degree double. We still need to show that these cycles generalize to arbitrary rooted neighborhoods, which is an important part of this result.
This also shows the reach that SSNC has across mathematics. Among the published literature there is not much work in the field of abstract algebra connected to the SSNC. However, given the importance of cycles in keeping nodes from having their degrees double, we are able to see that we are always able to find a representation that maps nodes to interior degrees in that many disjoint cycles. The disjoint cycle is what allows the interior degrees to double, because each cycle has nodes of degree one on the cycle itself.
By finding a positive answer to this localized version of the conjecture, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture. Not only that, but it helps us to dissect the problem down into two components: interior degree and exterior degree. If we know that the interior degree of every node inside the rooted neighborhoods doubles, then we can turn our attention to the exterior degrees.
We will then be able to introduce a traversal algorithm that explores these exterior degrees. It will systematically search the nodes of $G$ for a node whose exterior degree doubles. That is, we will question whose second (exterior) neighborhood is at least as large as its first. Lemma 6.6 (Decreasing Exteriors Lemma) will show that these exterior neighborhoods must shrink (decrease) at each step. We will then use Lemma 6.6 to show that we can generalize the interior load balancing lemma. Once we have proved these two major results, it will go a long way towards proving our contradiction.
The existence of a node whose degree doubles is underscored by Conjecture 3.1, which posits a different doubling requirement than Conjecture 1.1 that we have been working with. The Square Conjecture 3.1 asks for a node $u\in G^{2}$ such that $|N^{++}(u)|\geq 2\cdot|N^{+}(u)|$. We have equivalence between Conjecture 1.1 and Conjecture 3.1, so we can use both interchangeably. Once we show that the DNSP mandates the shrinking of exterior degrees, it will inevitably lead to a node whose exterior degree doubles. We will add to this Lemma 6.4 (Interior Degrees Double) which states that all node’s interior degrees double, and we arrive at a violation of the DNSP for that node. The combined doubling of both interior and exterior degrees for any single node would mean $|N^{++}(u)|\geq|N^{+}(u)|$, directly contradicting the DNSP and thus proving the Second Neighborhood Conjecture.
Lemma 6.5.
Exterior Degree Doubling Let $G$ be an oriented graph represented by a Graph Level Order with root node $v_{0}$. Further , assume $G$ has no back arcs. Suppose $u_{i}\in R_{i}$ has its interior out-degree double and $\text{ext}(u_{i-1},u_{i})=R_{i+1}$ Then if $|R_{i+1}|\geq|R_{i+2}|$, $u_{i}$ will have its exterior out-degree double as well, causing the overall out-degree of $u_{i}$ to double.
Proof.
Let us begin by restating our definitions. For a node $u_{i}$, we define the interior out-degree with respect to a parent $u_{i-1}$ as
$|\text{int}(u_{i-1},u_{i})|.$
Similarly, we define the exterior out-degree with respect to that same parent as
$|\text{ext}(u_{i-1},u_{i})|.$
By the assumption that $u_{i}$’s interior out-degree doubles, we are assuming that the number of second interior neighbors of $u_{i}$ with respect to $u_{i-1}$ is at least as large as its number of first interior neighbors. That is,
$|N^{++}_{\text{int}}(u_{i})|\geq|N^{+}_{\text{int}}(u_{i})|.$
By Lemma 5.1 (Partition of Node’s Degree), and the lack of back arcs in $G$, the behavior of a node’s degree doubling is dependent on its interior out-degree doubling and its exterior out-degree doubling. Because these sets are disjoint, they act independently of one another. Thus, if we can also show that $u_{i}$’s exterior out-degree doubles, it will show that $u_{i}$’s overall degree doubles.
By assumption, the nodes of $\text{ext}(u_{i-1},u_{i})$ are in $R_{i+1}$. These are the first exterior out-neighbors of $u_{i}$. The nodes in $R_{i+2}$ are the first out-neighbors of the nodes in $R_{i+1}$, thus they are the second exterior out-neighbors of $u_{i}$. The assumption that $|R_{i+1}|=|R_{i+2}|$ says that the node $u_{i}$ has an equal number of first exterior out-neighbors and second exterior out-neighbors. That is,
$|N^{++}_{\text{ext}}(u_{i})|=|N^{+}_{\text{ext}}(u_{i})|.$
This makes $u_{i}$ satisfy the degree doubling condition for its exterior degree. We have already shown that $u_{i}$ satisfies the degree doubling condition for its interior degree, so $u_{i}$ satisfies the degree doubling condition for its overall degree.
This completes the proof. ∎
Lemma 6.5 is the counterpart of Lemma 6.4. It will help us with our traversal algorithm, as we will walk through the graph and seek conditions that will make a node’s exterior degree double. This lemma formalizes that when we have two neighborhoods that are equal in size, there is a node whose degree doubles.
Lemma 6.6.
Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a Graph Level Order, where there are no back arcs, with $u_{i}\in R_{i}$ and
$v_{i+1}\in N^{+}(u_{i})\cap R_{i+1},$
where both $u_{i}$ and $v_{i+1}$ have the Decreasing Neighborhood Sequence Property (DNSP) and the node $v_{i+1}$ has its interior degree double. Then, for all
$z\in\text{ext}(u_{i},v_{i+1}),$
we have
$|\text{ext}(v_{i+1},z)|<|\text{ext}(u_{i},v_{i+1})|.$
That is, the exterior degree of $v_{i+1}$ with respect to $u_{i}$ decreases for every second neighbor $z$ of $u_{i}$.
Proof.
Let us first restate our definitions. The DNSP say that a node’s second neighborhood is strictly less than its first neighborhood
$|N^{++}(u)|<|N^{+}(u)|.$
Recall that for an arc $u_{i}\to v_{i+1}$,
$\text{int}(u_{i},v_{i+1})=N^{+}(x)\cap N^{+}(v_{i+1}).$
Similarly, for that same arc $u_{i}\to v_{i+1}$,
$\text{ext}(u_{i},v_{i+1})=N^{++}(u_{i})\cap N^{+}(v_{i+1}).$
And for the arc $v_{i+1}\to z$,
$\text{ext}(v_{i+1},z)=N^{++}(v_{i+1})\cap N^{+}(z).$
Assume that we have a node $v_{i+1}\in R_{i+1}$, whose interior degree doubles. We assumed there are no back arcs, which means that for all
$\forall w\in\text{ext}(u_{i},v_{i+1})\cap\text{ext}(v_{i+1},z)=\emptyset.$
For the sake of contradiction, suppose that there exists a $z\in\text{ext}(u_{i},v_{i+1})$ such that
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|.$
We can break the degree of $v_{i+1}$ down into two components (remember there are no back arcs that would be a third component), its interior degree and its exterior degree
$d^{+}(v_{i+1})=|\text{int}(u_{i},v_{i+1})|+|\text{ext}(u_{i},v_{i+1})|.$
By assumption, the node $v_{i+1}$ has its interior degree double. Thus, we have that if we can show that $v_{i+1}$’s exterior degree also doubles, it would lead to a contradiction because $v_{i+1}$ would have both its interior and exterior degree doubling, and thus its overall degree doubling, violating the DNSP.
Our assumption for the sake of contradiction was that $z\in\text{ext}(u_{i},v_{i+1})$ such that
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|.$
The node $v_{i+1}$’s out-degree will be impacted by $\text{ext}(u_{i},v_{i+1})$ and $\text{ext}(v_{i+1},z)$. This means that
$z\in\text{ext}(u_{i},v_{i+1})=N^{++}(x)\cap N^{+}(v_{i+1})=R_{i+2}$
$\begin{split}|\text{ext}(v_{i+1},z)|&\geq|\text{ext}(u_{i},v_{i+1})|\\ |\text{ext}(v_{i+1},z)|+|\text{ext}(u_{i},v_{i+1})|&\geq|\text{ext}(u_{i},v_{i+1})|+|\text{ext}(u_{i},v_{i+1})|\\ |\text{ext}(v_{i+1},z)|+|\text{ext}(u_{i},v_{i+1})|&\geq 2\cdot|\text{ext}(u_{i},v_{i+1})|\end{split}$ (1)
Thus we have the following conditional:
If
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|,$
then
$|\text{ext}(v_{i+1},z)|+|\text{ext}(u_{i},v_{i+1})|\geq 2\cdot|\text{ext}(u_{i},v_{i+1})|$
We have two exterior sets on the left of the equation that we would like to sum, but we need to be sure that there is no node
$w\in\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1}).$
Because this is a sum of set sizes, to correctly compare the sizes of these exterior sets, we first need to ensure that they are disjoint. We will define a set to represent their intersection and denote it $I$.
Consider the set
$I=N^{++}(v_{i+1})\cap N^{+}(z)\cap N^{++}(u_{i})\cap N^{+}(v_{i+1}).$
The set
$I(x,v_{i+1},z)=\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1})$
is the intersection of those two sets.
Our job is to determine the cardinality of
$I(u_{i},v_{i+1},z).$
Namely can there exist a
$w\in I(u_{i},v_{i+1},z).$
$w\in I$ implies that there are two possibilities for $w$. Either
$w\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$
or
$w\in N^{++}(v_{i+1})\cap N^{+}(z).$
If
$w\in N^{++}(u_{i})\cap N^{+}(v_{i+1}),$
This would mean that there was an arc $u_{i}\to w$ across two rooted neighborhoods, skipping all nodes in the rooted neighborhood $R_{i+1}$. We do not allow that through our shortest path constructions though.
If
$w\in N^{++}(v_{i+1})\cap N^{+}(z),$
then we have an arc from the rooted neighborhood $R_{i+2}$ where $z$ lies, to $R_{i+1}$ where $v_{i+1}$ and $w$ would be. This arc $z\to w$ would represent a back arc though, which we assumed did not exist in this graph. This would mean that the exterior degree of $v_{i+1}$ is at least as large as the exterior degree of $u_{i}$.
From $|\text{int}(u_{i},v_{i+1})|$ doubling, $v_{i+1}$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DNSP. The absence of back arcs prevents any increase in $v_{i+1}$’s exterior degree from its second neighbors.
Thus, the assumption
$|\text{ext}(v_{i+1},z)|\allowbreak\geq\allowbreak|\text{ext}(u_{i},v_{i+1})|$
leads to a contradiction, confirming that
$|\text{ext}(v_{i+1},z)|\allowbreak<\allowbreak|\text{ext}(u_{i},v_{i+1})|$
for all
$z\allowbreak\in\allowbreak\text{ext}(u_{i},v_{i+1}).$
Therefore, the lemma holds: for all $z\in\text{ext}(u_{i},v_{i+1})$, the exterior degree of $v_{i+1},z$ is strictly smaller than the exterior degree of $u_{i},v_{i+1}$. In other words, the size of the exterior neighbors decreases between consecutive neighborhoods. ∎
This lemma is important because, in contrast to other areas in Graph Level Order representations, we actually see at a set-theoretic level how back arcs can disrupt the Graph Level Order. It is not just dual paths, but it is in overlapping exterior sets. Once again, this must be addressed carefully.
Section 7 (Back Arcs) does address this as well. Even though this is a different issue. It is still resolved by the no back arcs lemma we present there. When back arcs are present, we will not be able to present the oriented graph in a Graph Level Order. Instead, they will still lead directly to a degree doubling node. We will still be able to take advantage of splitting up a nodes degree equation. The notion of interior degree doubling has not been impacted by back arcs and we will be able to take advantage of that. These back arcs will lead to second neighbors of the source node and a different path towards this doubling algorithm.
$u_{i}$$v_{i+1}$wz
Figure 11: Illustration of a back arc causing exterior neighbors to have overlap. The arc $u_{i}\to v_{i+1}$ has exterior neighbors $w\text{ and }z$. The arc $v_{i+1}\to z$ has one exterior neighbor, $w$. We see that the node $w$ is in both sets.
Figure 11 illustrates the situation that can cause the two exteriors on the left hand side in the proof in 6.6. The node $u_{i}$ has one out-neighbor, $v_{i+1}$. The node $v_{i+1}$ has two out-neighbors, $\{w,z\}$. This makes $\text{ext}(u_{i},v_{i+1})=\{w,z\}$. To calculate $\text{ext}(v_{i+1},z)$, we first calculate the nodes at distance two from $v_{i+1}$. Traversing through the back arc, $z\to w$. Then to traverse the node of distance one, we just traverse that same arc and see that $\text{ext}(v_{i+1},z)=\{w\}$.
More importantly though, what Lemma 6.6 (Decreasing Exteriors) showcases is that as we proceed on a path through exterior nodes from a minimum degree node $v_{0}$, the sizes of these exterior sets must decrease, as long as there are no back arcs along the path. We were first able to partition a node’s degree into interior and exterior degree by Lemma 5.1 (Partition of Node’s Degree). The point of Lemma 6.4 (Interior Degree Double) was to show that once this is done, the interior portion of this degree will always double. Now, this lemma and the DNSP are the foundations of our traversal algorithm through exterior neighbors.
For a more concrete example, consider the node $v_{0}$. This lemma is saying that the size of each exterior set we traverse, first through a node in $R_{1}$, then a node in $R_{2}$, etc, must be bounded first by $\delta-1$, then by $\delta-2$, continuing on throughout the graph. As we see, this is a much stronger lemma for graph traversal than Lemma 6.3 (Exterior Load Balancing).
Lemma 6.7.
Generalized Load Balance Lemma Suppose $G$ is an oriented graph with a node $v_{0}$ of minimum out-degree, and define rooted neighborhoods $R_{0},R_{1},\ldots,R_{k}$ based on distance from, constructed without back arcs. Assume $G$ is a minimal counterexample to the Second Neighborhood Conjecture (SSNC), i.e., every node $x\in G$ satisfies the Decreasing Neighborhood Sequence Property (DNSP): $|N^{++}(x)|<|N^{+}(x)|$.
Then, for any node $u_{i}\in R_{i}$ with $i\geq 1$, and any parent $u_{i-1}\in R_{i-1}$ of $u_{i}$, the number of common out-neighbors satisfies $|N^{+}(u_{i-1})\cap N^{+}(u_{i})|\geq i$.
Proof.
Let us first restate our definitions. The DNSP says that a node’s second neighborhood is strictly less than its first neighborhood $|N^{++}(u)|<|N^{+}(u)|$. For an arc $u\to v$ with $u\in R_{i-1}$ and $v\in R_{i}$, define: The interior out-degree of $v$ with respect to $u$ as
$|\text{int}(u,v)|:=|N^{+}(u)\cap N^{+}(v)|.$
The exterior out-degree of $v$ with respect to $u$ as
$|\text{ext}(u,v)|:=|N^{++}(u)\cap N^{+}(v)|.$
The interior neighbors will lie in the rooted neighborhoods $R_{i}$.
For any $v$, we are still undergoing the practice of splitting its total out-degree into its interior degree and its exterior degree. That is, its degree equation can be stated as $d^{+}(v)=|\text{int}(u,v)|+|\text{ext}(u,v)|$. We proceed by induction on distance $i$ from $v_{0}$:
Base case (i=1): Let $u_{1}\in R_{1}$ with parent $v_{0}\in R_{0}$. This is exactly what we saw in Lemma 6.2 (Interior Load Balancing), with $|N^{+}(v_{0})\cap N^{+}(u_{1})|\geq 1$ ,so the base case holds.
Induction hypothesis: Assume for all $1\leq k\leq i$, and any $u_{k}\in R_{k}$ with parent $u_{k-1}\in R_{k-1}.$
$|N^{+}(u_{k-1})\cap N^{+}(u_{k})|\geq k$
Inductive step: Consider any $u_{i+1}\in R_{i+1}$ with parent $u_{i}\in R_{i}$. Suppose, for contradiction,
$|N^{+}(u_{i})\cap N^{+}(u_{i+1})|<i+1$
i.e.,
$|\text{int}(u_{i},u_{i+1})|\leq i$
The minimum out-degree node $v_{0}$ says that the node $u_{i+1}$ must have an out-degree of at least $\delta$. Since $u_{i+1}$ has out-degree at least $\delta$, we have that
$|\text{ext}(u_{i},u_{i+1})|=d^{+}(u_{i+1})-|\text{int}(u_{i},u_{i+1})|\geq\delta-i.$
Lemma 6.6 (Decreasing Exteriors) states that because we have interior degree at least $k$ for every node $u_{k}\in R_{k}$ along the path from $v_{0}$ to $u_{i}$ (the parent node of $u_{i+1}$), exterior degrees must strictly decrease as we traverse outwards. Formally, for nodes $u_{k}\in R_{k}$ and $u_{k+1}\in R_{k+1}$, we have the following equation:
$|\text{ext}(u_{k},u_{k+1})|>|\text{ext}(u_{k+1},z)|$
for any $z\in\text{ext}(u_{k},u_{k+1})$.
This implies exterior degrees along the path decrease from $\delta$ down by at least 1 at each step, bounding
$|\text{ext}(u_{i},u_{i+1})|<\delta-i.$
But this contradicts the earlier inequality that
$|\text{ext}(u_{i},u_{i+1})|\geq\delta-i.$
This contradiction arises from assuming the interior degree is less than $i$+1. Hence, the assumption is false, and we conclude:
$|N^{+}(u_{i})\cap N^{+}(u_{i+1})|\geq i+1$
This completes the induction and proves the lemma. ∎
The Lemma 6.7 (Generalized Load Balance) is a generalization of 6.2 (Interior Load Balancing Lemma). Now we see that in order for a node $u$ in a rooted neighborhood $R_{i}$ to not have its degree double, we need all its out-neighbors $N^{+}(u)$ to have their interior out-degree set to at least $i$. This is saying that each node in the rooted neighborhood $R_{i}$ must relate to at least $i$ other nodes in that same rooted neighborhood.
As mentioned above, our initial lemma 6.2 exists in literature on the Seymour Conjecture [6]. The truly critical insight lies in the generalization of the load-balancing principle we have uncovered within the Graph Level Order.
The inductive step of our reasoning reveals that the prevention of degree doubling is not a phenomenon limited to specific initial conditions but rather a fundamental requirement operating throughout the entire hierarchical structure.
Every node, at every level of the Graph Level Order, must participate in effectively distributing outgoing connections to the subsequent rooted neighborhood. This collective load balancing is essential to ensure that predecessors at earlier levels do not experience a doubling of their degree. The failure of any node within any rooted neighborhood to adequately contribute to this distribution inevitably leads to degree doubling of a predecessor.
This generalized principle underscores the intricate and interconnected nature of the Graph Level Order, highlighting that the avoidance of degree doubling is a global constraint dependent on the cooperative behavior of nodes across all rooted neighborhoods, far beyond the easily observable base cases.
Moreover, the inductive step made use of Lemma 6.4 (Interior Degree Doubling) and Lemma 6.6 (Decreasing Exteriors) to reach a contradiction. Now that we have an assumption of each node needing to support, not only their parents but their ancestors as well through higher degrees in further rooted neighborhoods, we can see the usage of the constructive 6.4 mapping algorithm. We know that there will always exist feasible cycle mappings for these nodes, no matter how large the rooted neighborhood. And because we know cycles immediately lead to degree-doubling nodes, this solves the interior degree case in $O(1)$ complexity. What remains is the traversal algorithm through the exterior nodes and back arcs.
Figure 12: This figure illustrates the Generalized Load Balance Lemma, showing how the required interior degree increases as the distance from $v_{0}$ increases, while the number of exterior arcs decreases.
Node $R_{1}$ Targets $R_{2}$ Targets $R_{3}$ Targets Neighborhood
0 1, 2, 3, 4, 5, 6 $R_{0}$
1 2, 3, 4 8, 9, 10 $R_{1}$
2 3 7, 8, 9, 10, 11 $R_{1}$
3 4 7, 8, 9, 10, 11 $R_{1}$
4 5 7, 8, 9, 10, 11 $R_{2}$
5 6 7, 8, 9, 10, 11 $R_{2}$
6 1 7, 8, 9, 10, 11 $R_{3}$
7 8 13, 14, 15, 16 $R_{3}$
8 9 13, 14, 15, 16 $R_{3}$
9 10 13, 14, 15, 16 $R_{3}$
10 11 13, 14, 15, 16 $R_{3}$
11 8 13, 14, 15, 16 $R_{3}$
Table 3: A JSON representation (in table form) of a Graph Level Order with an irregular $R_{1}$ interior degree but exhibiting load balancing. The interior degrees in $R_{1}$ vary (irregularity), yet sufficient internal connections within $R_{1}$ maintain overall balance, preventing degree doubling in $R_{0}$.
We have shown so many examples with regular interior degrees that one can grow to expect that as the only possibility. The Example 3 is a situation where the minimum degree node $v_{0}$ has degree 6. Then in the first rooted neighborhood we have the node 1 that has two interior neighbors, $2,3$. All other nodes in that rooted neighborhood only have one interior neighbor in $R_{1}$. This is not a problem for the Decreasing Neighborhood Sequence Property though. In fact, a node relating to more nodes in an earlier container implies that it will either (1) relate to less nodes in a later container, thereby still complying with Lemma 6.6 (Decreasing Exteriors). The alternative is that a node that has a higher interior degree could have a higher overall degree. This node would still be bound though by 6.6 (Decreasing Exteriors), to prevent its parents from having their degrees double. The fact that all nodes in the first rooted neighborhood still have interior degree of at least 1 means that they are all participating in a cycle, which means that they are all contributing to the load balancing.
What we will see through the next lemma is that we can in fact use this lemma to bound the size of entire rooted neighborhoods.
Lemma 6.8.
DNSP Impact on Rooted Neighborhood Size Let $G$ be an oriented graph rooted at a minimum out-degree node $v_{0}$, with rooted neighborhoods $R_{0},R_{1},\dots,R_{k}$ constructed by layering (i.e., $R_{i}$ is the set of nodes at distance $i$ from $v_{0}$). Suppose all nodes satisfy the Decreasing Neighborhood Sequence Property (DNSP), and that the graph contains no back arcs. Then for all $i\geq 0$,
|$|\{y\in\text{ext}(u,v)\text{ s.t. }u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Proof.
We proceed by induction on the distance $i$.
Base Case ($i=0$): The neighborhood $R_{0}=\{v_{0}\}$, and its out-neighbors form $R_{1}$, with $|R_{1}|=d^{+}(v_{0})$ by definition. Now consider $R_{2}$, which consists of nodes reachable by an arc from nodes in $R_{1}$. We assume that all nodes, including $v_{0}$ and all nodes in $R_{1}$ have the DNSP. Let $v_{0}\in R_{0}$, $v\in N^{+}(v_{0})=R_{1}$, and define $\text{ext}(v_{0},v):=N^{++}(v_{0})\cap N^{+}(v)$. From DNSP holding for $v_{0}$, we have
$|N^{++}(v_{0})|<|N^{+}(v_{0})|=d^{+}(v_{0}),$
So no matter what set we intersect $N^{++}(v_{0})$ with from $R_{1}$, the resulting intersection $N^{++}(v_{0})\cap N^{+}(v)$ will have a smaller cardinality than $d^{+}(v_{0}$.
This is the bound we need for the exterior set of $v_{0}\to v$.
$|\text{ext}(v_{0},v)|\leq d^{+}(v_{0})-1.$
Inductive Hypothesis: Assume that for some $i\geq 0$,
|$|\{y\in\text{ext}(u,v)\text{ s.t. }u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Inductive Step: Consider $u\in R_{i+1}$, and $u\to v\in G$, where $v\in R_{i+2}$, and we will examine $\text{ext}(u,v):=N^{++}(u)\cap N^{+}(v)$.
From Lemma 6.6 (Decreasing Exteriors Lemma), we have:
$|\text{ext}(u,v)|<|\text{ext}(w,u)|,$
for any parent $w\in R_{i}$, $u\in R_{i+1}$.
By the inductive hypothesis,
$|\text{ext}(w,u)|\leq d^{+}(v_{0})-i.$
We can combine these inequalities to get,
$|\text{ext}(u,v)|<d^{+}(v_{0})-i\to|\text{ext}(u,v)|\leq d^{+}(v_{0})-(i+1).$
Hence, the bound holds for level $i+1$. This completes the induction and proves the lemma. ∎
What this has established through the DNSP, and contingent upon the proof of the generalization of load balancing, is that the rooted neighborhoods are decreasing in size as we move further away from the minimum out-degree node $v_{0}$. We spoke of the DNSP being a bottleneck, and here is another example where, across the entire graph, we are actually seeing it restrict the sizes of rooted neighborhoods. DNSP is a local bottleneck that induces a global contraction of the neighborhood layers in the Graph Level Order.
We began with the assumption that the Decreasing Neighborhood Sequence Property (DNSP) established that for a node $u$, the size of its first out-neighborhood is strictly larger than its second out-neighborhood $|N^{+}(u)|>|N^{++}(u)|$. We have now witnessed the emergence of this principle at the level of entire rooted neighborhoods within the Graph Level Order structure. Lemma 6.8 (DNSP Impact on Rooted Neighborhood Size) shows that consecutive neighborhoods must strictly decrease in size $(|R_{i+1}|<|R_{i}|$ for $i\geq 1)$. This demonstrates DNSP rising from a local node-level constraint to a global feature of the graph’s partitioning. This highlights the penetrating and fundamental nature of the DNSP. It is felt at every node of the graph, impacting not only the individual nodes but the hierarchies above them as well. Because the Graph Level Order is organized in this way, it merges the power of the DNSP to truly organize these levels. This rooted neighborhood-level constraint of the DNSP sets the stage for the next level of understanding between the implications for the existence of back arcs. These back arcs hold weight in whether the Graph Level Order is just another data structure or if it can help solve the SSNC.
We have shown $|R_{i+1}|<|R_{i}|$, which means $|R_{i+1}|\leq|R_{i}|-1$. Now that we know that the rooted neighborhoods decrease in size, we can quantify the bounds on these sizes based on the distances from the minimum out-degree node. This will help with factors like back arcs and the ability to generalize the above lemma.
Lemma 6.9.
Formula for Rooted Neighborhood Size Let $G$ be an oriented graph with minimum out-degree node $v_{0}$, where the out-degree of $v_{0}$ is $\delta$. Suppose all nodes in $G$ satisfy the DNSP. Then the rooted neighborhoods $R_{i}$ satisfy the bounds:
$|R_{0}|=1,$
$|R_{1}|=\delta,$
$|R_{i}|\leq\delta-(i-1)$
Proof.
We proceed by induction on the distance $i$ of rooted neighborhoods $R_{i}$.
Base cases:
  • $R_{0}=\{v_{0}\}$, so $|R_{0}|=1$ by definition.
  • $R_{1}=N^{+}(v_{0})$, so $|R_{1}|=\delta$, again by definition.
The DNSP says $|N^{+}(v_{0})|>|N^{++}(v_{0})|$. All of the out-neighbors of $v_{0}$ are in the rooted neighborhood $R_{1}$ and every second neighbor of $v_{0}$ will be in the rooted neighborhood $R_{2}$. By Lemma 6.2 (Interior Load Balancing), these children $u\in R_{1}$ help prevent $v_{0}$ from becoming a degree doubling node by forming cycle(s) in $R_{1}$. This bounds the number of external neighbors in $R_{2}$ and thus the size of $|R_{2}|$.
Inductive hypothesis: Assume that for some $k<i$, we have:
$|R_{k}|\leq\delta-(k-1)$
Inductive step: We consider the neighborhood $R_{k+1}$. We apply Lemma 6.8 (DNSP Impact on Neighborhood Size), which implies:
$|R_{k+1}|<|R_{k}|\leq\delta-(k-1)\to|R_{k+1}|\leq\delta-((k+1))-1)=\delta-k.$
We conclude that the inductive structure holds for all $i$, giving:
$|R_{i}|\leq\delta-(i-1),$
What we saw from 6.8 (DNSP Impact on Rooted Neighborhood Size) is that the rooted neighborhoods decreased in size. This decrease must be at least by one at each neighborhood. What 6.9 (Formula for Rooted Neighborhood Size) does is combine this into a worse case calculation size for each rooted neighborhood. In addition, this allows us to calculate the worst case sizes for interior and exterior neighborhood sizes. This gives us a way to evaluate a node’s second neighborhood against these rooted neighborhood sizes, as a comparison for degree-doubling nodes.
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample. We also have an understanding of the Decreasing Neighborhood Sequence Property which will guide the nodes throughout the rest of this proof process.
What we see from the generalizations of the load balancing lemmas is that Lemmas 6.2 (Interior Load Balance) and 6.3 (Exterior Load Balance) were not just exceptions, they are the key building blocks of an essential structure. With these generalizations, we have both lower bounds on the necessary interior degrees and upper bounds on the exterior degrees required to keep a node’s degree from doubling. This is key for the load balance portion.
What we also saw in this section, though, is that as this load is passed from parent to child the size of these neighborhoods is reduced. This leads to a sequence of rooted neighborhoods that are getting smaller and smaller, but these smaller neighborhoods are expected to hold more nodes inside them. Obviously there is a collision that is inevitable. We need to ensure that other concepts like back arcs do not get in the way. The Decreasing Neighborhood Sequence Property (DNSP) says that we should continuously have these rooted neighborhoods decrease forever. Any time that they stop decreasing, we will arrive at a node, or set of nodes, whose degree doubles. This would be a violation of the DNSP assumption and a contradiction.
7 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate the influence these back arcs have on our data structure and our search for a counterexample. Understanding back arcs is crucial because their presence can invalidate certain structural assumptions that underpin our search for a counterexample, including violating the principles of the partition. Their presence can create shortcuts in the graph, potentially leading to a node’s second out-neighbor overlapping with earlier rooted neighborhoods in ways that violate the distance-based ordering of the Graph Level Order.
It is important to note that the back arcs and the exterior out-neighbors are distinct concepts. Back arcs are not a subset of exterior out-neighbors. While back arcs may connect to any previous rooted neighborhood, exterior out-neighbors are limited to connecting with the next rooted neighborhood. In addition, exterior out-neighbors do not represent all possible future rooted neighborhoods. Rather, an exterior out-neighbor $\text{ext}(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$. However, this is mainly a conceptual thing. The fact that back arcs do not point in a forward direction and exterior arcs do, does not change the fact that both exterior arcs and back arcs point a parent node to a node in a different rooted neighborhood. Therefore, it is both reasonable and useful to treat back arcs as a subclass of exterior arcs. They should be distinguished not by their directional flow but by the fact that they cross boundaries of the rooted neighborhoods.
If we look at Example 4.2 again, remember that this example was important because of the cycle among $v_{1},v_{2},v_{3}$. We see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|\text{int}(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $v_{1}\to v_{2}\in G$. Similarly, we see that $|\text{ext}(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $v_{1}\to v_{4}\in G$ and $v_{1}\to v_{5}\in G$ and $v_{0}\to v_{4}\in G$ and $v_{0}\to v_{5}\in G$.
Something we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. The next lemma formalizes that concept.
7.1 Consequences of Back Arcs
The key assumption of the Graph Level Order is that the rooted neighborhoods are disjoint. While the rooted neighborhoods form a partition of the nodes of the graph— ensuring that each node belongs to exactly one rooted neighborhood—arcs will respect this partition. This distinction becomes very important. In theory, we do not want arcs that are able to influence the nature of the data structure. However, back arcs—edges that go from a higher-indexed neighborhood back to a lower one—can cause overlap between exterior sets. These arcs introduce connections that undermine the neat separation the rooted neighborhoods imply. As such, the validity of the lemma depends not just on the rooted neighborhood structure but on the absence or control of these back arcs.
Recall that a back arc from a node in a rooted neighborhood $R_{i}$ is an arc directed to a node in neighborhood $R_{j}$, where $j<i$. These back arcs introduce a non-empty intersection between the sets of nodes reachable from $R_{i}$ and $R_{j}$. In the absence of back arcs, these rooted neighborhoods (or, more precisely, the sets of nodes relevant to our arguments) are disjoint. Therefore, we must carefully examine how the introduction of back arcs disrupts the disjoint nature and impacts our proofs. The rooted neighborhoods are members of a partition and thus are disjoint by definition.
Lemma 7.1.
Back Arcs Necessary Let $G$ be an oriented graph with minimum out-degree $v_{0}$ and node rooted neighborhoods $R_{i}$ and $R_{j}$, with $i>j$. If there are no back arcs from any node in $R_{i}$ to any node in $R_{j}$, then $R_{i}$ and $R_{j}$ are disjoint (i.e., $R_{i}\cap R_{j}=\emptyset$).
Proof.
Assume, for the sake of contradiction, that $R_{i}$ and $R_{j}$ are not disjoint. Then there exists a node $v$ such that $v\in R_{i}\cap R_{j}$. This means that $v$ is in both $R_{i}$ and $R_{j}$. By the definition of $R_{i}$, this means that $v$ is at distance $i$ from $v_{0}$. Likewise, by the definition of $R_{j}$, this means that $v$ is at distance $j$ from $v_{0}$. We assumed that $i\neq j$, so we are saying that a node $v$ is at two different distances from $v_{0}$. This is a contradiction, as a node cannot simultaneously be at distance $i$ and $j$ from $v_{0}$ when $i\neq j$, because there is a unique distance from $v_{0}$ to any node. Therefore, $R_{i}$ and $R_{j}$ must be disjoint. In other words, when there are no back arcs, then different rooted neighborhoods are disjoint. ∎
Lemma 7.1 sets up a situation where we first establish the presence of the Graph Level Order. We want to establish that an oriented graph without back arcs will have no intersecting rooted neighborhoods. Our proof proceeds by contradiction, supposing that a node is in two different rooted neighborhoods. The problem with this assumption is that the definition of a node belonging to a rooted neighborhood $R_{i}$ means that node is distance $i$ from the minimum out-degree node $v_{0}$. So a node $v$ being in two different rooted neighborhoods means it is at two different distances from $v_{0}$. This is not possible, though, and we conclude that they must be disjoint.
Lemma 7.2.
Back Arcs Sufficient Let $G$ be an oriented graph with minimum degree node $v_{0}$ and rooted neighborhoods $R_{i}$ and $R_{j}$, where $j<i$. If $R_{i}$ and $R_{j}$ are not disjoint, then there exists a back arc from a node in $R_{i}$ to a node in $R_{j}$.
Proof.
Suppose that $R_{i}$ and $R_{j}$ are not disjoint. Then there exists a node $v$ such that $v\in R_{i}\cap R_{j}$. This means that $v\in R_{i}$ and $v\in R_{j}$. Since $i<j$, and both $v\in R_{i}$ and $v\in R_{j}$, there must be a path from $v_{0}$ to $v$ of length $i$ and a path from $v_{0}$ to $v$ of length $j$. The only way this is possible is if there is a back arc from a node in $R_{i}$ to a node in $R_{j}$. Then we would extend the path from $v_{0}\to R_{i}\to v_{1}\to back\_arc\to R_{j}\to v_{2}$. Thus, the existence of non-disjoint rooted neighborhoods, when $i<j$, implies the existence of a back arc.
If there were no back arc from $R_{i}$ to $R_{j}$, then any path from $v_{0}$ to a node in $R_{i}$ would have length $i$, and any path to a node in $R_{j}$ would have length $j$. If a single node $v$ were in both $R_{i}$ and $R_{j}$, it would imply two different shortest path lengths from $v_{0}$ to $v$, which is impossible in a simple directed graph. ∎
Lemma 7.2 (Back Arcs Sufficient) sets up the necessary conditions in which two neighborhoods can intersect. This lemma is simply a contrapositive of Lemma 7.1 (Back Arcs Required). We can see a visualization of this in Example 7.1. We see a situation where we have five (or more) arcs, $y\to w$, $x\to y\to\ldots\to u\to z\to w$.
Example 7.1.
[…]
In Example 7.1, we observe that the node $w$ is distance two from $x$ and distance one from $y$, i.e., in $\text{ext}(x,y)$. The nodes preceding the back arc are $u$ and $z$, so $\text{ext}(u,z)=\{w\}$. Without the arc $z\to w$, the set $\text{ext}(x,y)=\emptyset$. There may be other nodes on the path from $y$ to $u$, and we can calculate those exteriors as well. The truth about any such path is that its distance to $w$ is greater than two, so $w$ would not be in the set of exterior out-neighbors. The other possibility is that a node along the path from $y$ to $u$, is in the set $\text{ext}(x,y)$. This is not possible, though,because the element that will be in $\text{ext}(x,y)$ is the first element along the path, and that cannot be the element that is in the exterior out-neighbor from the path, as it has to be distance two from this first node.
7.2 Dealing With Back Arcs
The definition of a back arc is an edge $y\to z$, where $y\in R_{i}$, $z\in R_{j}$, and $j<i$, respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the rooted neighborhood organization, which contrasts with this.
Lemma 7.3 (No Back Arc) effectively outlines the fundamental problem with back arcs. Lemma 6.8 (DNSP Impact on Neighborhood Size) introduced how the DNSP impacted rooted neighborhoods, decreasing in size as we went further from the minimum out-degree node $v_{0}$. Then we were able to prove Lemma 6.7 (Generalized Load Balance), which showed that as we moved further from the minimum out-degree node, these interior nodes were still keeping up their part of the load balance. Finally, we were able to combine these two lemmas with Lemma 6.4 (Interior Degree Doubling), which made each rooted neighborhood serve as its own SSNP and inside the $R_{i}$, we find that the interior degree would double. When we combine these three lemmas, we see that the back arcs lead to a node whose rooted neighborhood size is always greater than the head node of the back arc’s exterior neighborhood’s size.
Lemma 7.3.
No Back Arcs Let $G$ be an oriented graph with rooted neighborhoods $R_{i}$ that are well-ordered by their indices. If a node $u_{i}\in R_{i}$ whose interior degree doubles neighborhood in $R_{i}$, and it has its first back arc to a node $v_{k}$ in neighborhood $R_{k}$, where $k<i$, meaning there is no back arc from $u_{i}$ to $R_{j}$ where $k<j<i$, then the size of the second out-neighbors $N^{++}(u_{i})$ is at least $\text{deg}(v_{k})$, and this will cause the total degree of $u_{i}$ to double as well.
Proof.
We assume here that all nodes in $G$ have the DNSP. Assume that $u_{i}\in R_{i}$ has its earliest back arc to $v_{k}\in R_{k}$ where $k<i$. A back arc $u_{i}\to v_{k}$ means $v_{k}$ is a first out-neighbor of $u_{i}$. The neighbors of $v_{k}$ are then at distance two from $u_{i}$ through $v_{k}$ (a path of length two: $u_{i}\to v_{k}\to N^{+}(v_{k})$).
By ’first back arc,’ we mean the back arc originating from $u_{i}$ that targets the rooted neighborhood with the largest index $k$ such that $k<i$ and there are no back arcs from $u_{i}$ to any $R_{j}$ where $k<j<i$.
We also have that by the interior degree doubling of $u_{i}$ and the lack of back arcs before $i$, we know that $u_{i}$ will have only interior and exterior arcs. By $u_{i}$ having its interior degree double, if we can show that its exterior degree will double, then its overall degree will double. We just showed that $w\in N^{+}(v_{k})$ was a second out-neighbor of $u_{i}$ through an exterior arc. By Lemma 6.9 (DNSP Impact on Neighborhood Size), we have that $\text{ext}(u_{i-1},u_{i})<\delta-i$, for some parent $u_{i-1}\in R_{i-1}$ of $u_{i}$. Similarly, because $v_{0}$ is a minimum out-degree node in $G$ we have that $d^{+}(v_{k})=|N^{+}(v_{k})|\geq\delta$. We can combine these two inequalities and we see that
$|N^{+}(v_{k})|\geq\delta>\delta-i>\text{ext}(u_{i-1},u_{i}).$
Simplifying, we see that
$|N^{+}(v_{k})|>\text{ext}(u_{i-1},u_{i}).$
Even the situation where $v_{k}\in R_{i-1}$ would not stop $u_{i}$ from becoming a degree-doubling node because Lemma 6.7 (Generalized Load Balance) states that $i$ of $u_{i}$’s neighbors must be interior neighbors. This causes the inequality to hold.
We need to consider the scenario that $u_{i}$ sends more back arcs to other nodes later than $v_{k}$. In this situation, we would have a similar calculation where the number of neighbors of this new back arc would be greater than the number of exterior arcs. Indeed, the inequities are even greater because these back arcs would bring in new nodes, and the exterior arcs stay the same for each back arc.
This causes the total degree of $u_{i}$ to double because $u_{i}$ has more second out-neighbors. Therefore, the back arc from $u_{i}$ to $v_{k}$ causes the second out-neighbors of $u_{i}$ to increase to at least $\text{deg}(v_{k})$, which is greater than the number of first out-neighbors of $u_{i}$, causing its degree to double. ∎
The proof above demonstrates that a back arc to an early neighborhood would simply lead to a degree-doubling node. Our main assumption here is that the first back arc from $u_{i}$ is to the node $v_{k}$. What is essential is the Graph Level Order that allows us to organize, not only the nodes themselves, but also the back arcs in a schematic way so as to be able to select the earliest one from $u_{i}$. Similarly, this data structure has been allowing us to define arcs by their relationship of their nodes to the distances of their rooted neighborhoods. Nodes with the same distances and the same parent are interior neighbors, nodes with a difference of 1 and the same parent are exterior neighbors, and nodes with a difference greater than 1 and the same parent are called back arcs. What the proof shows is that $v_{k}$ has a higher degree than $u_{i}$’s number of exterior neighbors, which is what causes $u_{i}$’s degree to double.
Once we have the DNSP, the logic for back arcs is simple: it sends flow from a node in a smaller rooted neighborhood to a larger rooted neighborhood. Previously, the lemmas in that section were held back because of the possibility of back arcs. However, we see now that back arcs do not hinder the DNSP logic at all. In fact, these arcs only help us find degree-doubling nodes.
8 Main Theorem
8.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some results.
So far, this battle has been a battle between a catalyst, $v_{0}$, and the bottleneck, the Decreasing Neighborhood Sequence Property (DNSP). We have proved that: Back arcs cannot exist; interior degrees must double inside each rooted neighborhood, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
8.2 Theorem Statement
Degree-Doubling by Exterior Decreasing
Theorem 8.1.
Let $G$ be an oriented graph where all nodes satisfy the Decreasing Neighborhood Sequence Property (DNSP). Let $v_{0}$ be a minimum out-degree node used to set up our rooted neighborhood partition. Then Seymour’s conjecture holds true for $G$.
Proof.
By Lemma 6.8 (DNSP Neighborhood Size Property), the sizes of rooted neighborhoods $|R_{1}|,|R_{2}|,\ldots$ form a strictly decreasing sequence of positive integers. Consequently, there exists a smallest $i>0$ such that
$|R_{i}|\leq 2.$
Let $u_{i-1}\in R_{i-1}$ be a parent of the nodes in $R_{i}$.
Case 1: If $|R_{i}|=1$, say $R_{i}=\{v\}$.
Then $u_{i-1}$ has an exterior out-neighbor $v$. Since $|R_{i}|$ is minimal, $|R_{i-1}|>1$. By Lemma 6.7 (General Load Balancing), $v$ should have an interior degree of at least $i$. However, with only one node in $R_{i}$, this is impossible without violating the DNSP at some previous node, which leads to degree doubling.
Case 2: If $|R_{i}|=2$, say $R_{i}=\{v_{1},v_{2}\}$.
Then the parent $u_{i-1}$ had at least two exterior neighbors. For $v_{1}$ and $v_{2}$ to maintain the load balancing, they would need interior connections. However, with only two nodes, at least one node must have all its outgoing arcs to $R_{i+1}$ or back to earlier neighborhoods, forcing a degree-doubling scenario for its predecessor in $u_{i-1}$. ∎
Theorem 8.2 (Degree-Doubling by Exterior Decreasing) is exactly what we saw in Section 4 (Initial Lemmas), in particular in Lemma 4.3. If we now view those as rooted neighborhoods, we can say we had the following rooted neighborhoods
Rooted Neighborhood Nodes Size
$R_{0}$ 0 1
$R_{1}$ 1, 2, 3 3
$R_{2}$ 4, 5 2
Table 4: A look at Lemma 4.3 through the lens of rooted neighborhoods and their sizes
What we see from this table is that once the size of $R_{2}$ reached only 2 nodes, we were guaranteed that some node in $R_{1}$ would have its degree double by the combination of size of $R_{2}$ and the interior degrees doubling. The nodes in $R_{1}$ have interior degrees of 1 and thus automatically have their degrees double in that neighborhood. The traversal algorithm would simply search for a node whose degree is less than 2. This is achieved in $R_{2}$ where at least one node must have the exterior degree of 3. This implies that all nodes in $R_{1}$ that are parents of this node will have their degrees double.
Corollary 8.1.
Degree-Doubling by Rooted Neighborhood Density In an oriented graph $G$ with a minimum out-degree node $v_{0}$ and rooted neighborhoods $R_{1},\allowbreak R_{2},\allowbreak\ldots\allowbreak R_{n}$, the shrinkage of neighborhoods and growth of interior degrees guarantees an inevitable collision.
Two scenarios emerge:
  • Case 1: The rooted neighborhoods shrink rapidly while interior degrees grow.
    Eventually, no oriented graph can support the required structure, forcing degree doubling.
  • Case 2: For very small minimum degree $\delta$, the neighborhoods shrink slowly and interior degree growth does not immediately force a collision before nodes run out.
Corollary 8.1 (Degree-Doubling by Rooted Neighborhood Density) captures the essence of shrinking rooted neighborhood sizes and an increased number of nodes. This was represented by the yin and yang effects early on. We see it come full circle here. Ultimately, the interior out-neighbors represented the number of interior nodes, while the exterior out-neighbors represented the size of the rooted neighborhoods.
In small graphs, such as the one shown in Example 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed. This is because the rooted neighborhood sizes decrease linearly (at a rate of $O(n)$), while the interior degrees increase linearly (also at a rate of $O(n)$). However, the number of arcs within the oriented graph grows quadratically, at a rate of $O(n^{2})$. Consequently, while the number of nodes increases linearly, the number of edges increases quadratically. In small graphs, the growth hasn’t progressed far enough for the quadratic edge growth to outpace the linear node growth and cause collisions. In larger graphs, however, this difference in growth rates becomes significant, and collisions begin to occur. This divergence implies that collisions must occur in large graphs, even if they’re absent in small cases.
Corollary 8.2.
Occurrence of Last Dense Rooted Neighborhood Let $G$ be an oriented graph with minimum degree $k$. Then there will be a rooted neighborhood that reaches maximal density.
Proof.
We begin by stating Lemma 6.7 (Generalized Load Balance). This states that as we move further from the minimum out-degree node, $v_{0}$, the nodes are forced to take on more of the load.
In particular, we see that in the rooted neighborhood $R_{i}$, each node must have an interior out-degree of at least $i$ in order to keep all of its predecessors in the chain of rooted neighborhoods from $v_{0}$ from having their degrees double.
Thus, we have a series of rooted neighborhoods that are getting smaller as we get further from the minimum out-degree node. At the same time, these rooted neighborhoods are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively leading to more densely rooted neighborhoods. As described in Corollary 8.1 (Degree-Doubling by Rooted Neighborhood Density), this cannot continue forever. There must be a last, most dense, rooted neighborhood. ∎
The SSNC is made more aesthetic by the appearance of these orientations of graphs as rooted neighborhoods get smaller. Not only has the conjecture been proven true, but with an inherent symmetry and order. There is a visual pattern emerging inside these rooted neighborhoods that is not just important to graph theorists but also possesses a structure that is easily appealing to non-mathematicians. Eventually, the constraint of increasing degree within shrinking rooted neighborhoods yields a unique rooted neighborhood where the number of arcs is maximized relative to the number of nodes—the densest rooted neighborhood.
Corollary 8.3.
Multiple Degree-Doubling Nodes Let $G$ be an oriented graph with minimum degree $\delta$, and let the most dense graph be in the rooted neighborhood $R_{i}$. Then every node in the rooted neighborhood $R_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one of abundance. There were some who doubted that one even existed. Now there are many.
Second is the way in which we find it. It is the presence of a last dense, rooted neighborhood with an oriented graph as these rooted neighborhoods decrease in size. We see this occur as the interior degrees are forced to increase as the distance from $v_{0}$ is increased. These interior degrees are set up to prevent a predecessor node’s degree from doubling. Ultimately, though, they are unable to do that, as we see the collision of the rooted neighborhood shrinkage and interior degree requirements.
Then third, we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group. Dense graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory. The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one of abundance.
8.3 Algorithm
Before we formally present the algorithm we have described in this paper, we will give two additional lemmas that will help with both the algorithm and the complexity of the algorithm. This algorithm leverages the Graph Level Order and the properties derived from the assumption of the Decreasing Neighborhood Sequence Property (DNSP) to detect a degree-doubling node. It systematically checks for back arcs, violations of the interior degree requirements, and the eventual shrinkage of rooted neighborhoods to a size that forces degree doubling in a predecessor.
Lemma 8.1.
One Interior Fail Means All Fail
Let $u_{i}\in R_{i}$ be a node such that for some parent $u_{p_{1}}\in R_{i-1}$, the interior degree condition fails:
$|\text{int}(u_{p1},u_{i})|<i.$
Then the same condition fails for every other parent $u_{p_{2}}\in R_{i-1}$, implying that every node in $R_{i-1}$ is a degree-doubling node.
Proof.
Suppose, for contradiction, there exists a node $u_{i}\in R_{i}$ and a parent $u_{p_{1}}\in R_{i-1}$ with
$|\text{int}(u_{p_{1}},u_{i})|<i).$
This implies that $u_{i}$ sends more than $\delta-i$ arcs to nodes in the next neighborhood $R_{i+1}$, i.e.,
$|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1.$
Consider any other parent $u_{p_{2}}\in R_{i-1}$ of $u_{i}$. Since arcs are oriented and back arcs are excluded, the neighbors of $u_{i}$ in $R_{i+1}$ are second out-neighbors of $u_{p2}$. Thus,
$|\text{ext}(u_{i},R_{i+1})|\leq|N^{++}(u_{p_{2}})|.$
But since
$|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1,$
and since $u_{p_{2}}$ has at most $\delta-i$ first out-neighbors (in $R_{i}$), we have
$|N^{++}(u_{p_{2}})|\geq\delta-i+1>\delta-i\geq|N^{+}(u_{p_{2}})|.$
Therefore,
$|N^{++}(u_{p_{2}})|\geq|N^{+}(u_{p_{2}})|,$
which means $u_{p_{2}}$ is a degree-doubling node, violating the DNSP.
Since this argument holds for $u_{p_{2}}\in R_{i-1}$, all parents of $u_{i}$ fail the interior degree condition, and thus all become degree-doubling nodes.
Lemma 8.2.
One Interior Succeed Means All Succeed Suppose that for a node $u_{p_{1}}\in R_{i-1}$ and every $u_{i}\in R_{i}$ the interior degree condition holds:
$|\text{int}(u_{p_{1}},u_{i})|\geq i.$
Then this condition holds for all other parents $u_{p_{2}}\in R_{i-1}$, and no node in $R_{i-1}$ is degree doubling.
Proof.
Assume that for a particular parent node $u_{p_{1}}\in R_{i-1}$, the interior arc condition holds for every child node $u_{i}\in R_{i}$:
$|\text{int}(u_{p1},u_{i})|\geq i.$
By definition, this implies that each such node $u_{i}$ sends at most $\delta-i$ arcs to the next neighborhood $R_{i+1}$, i.e.,
$|\text{ext}(u_{i},R_{i+1})|<\delta-i.$
Now, for any other parent $u_{p_{2}}\in R_{i-1}$, We want to show that for every shared child $u_{i}$, the interior condition also holds for $u_{p_{2}}$.
$N^{++}(u_{p_{2}})\subseteq\bigcup_{u_{i}\in N^{+}(u_{p_{2}})}\text{ext}(u_{i},R_{i+1}).$
Since each $u_{i}$ sends at most $\delta-i$ arcs to $R_{i+1}$, and by Lemma 6.9, the size of $R_{i+1}$ is also bounded by $\delta-i$, it follows that
$|N^{++}(u_{p2})|\leq|N^{+}(u_{p_{2}})|\times(\delta-i)\leq|N^{+}(u_{p_{2}}|\times 1,$
where the factor 1 applies if the out-degree matches the bound tightly. More precisely, the cardinality of second out-neighbors does not exceed that of first out-neighbors.
Hence,
$|N^{++}(u_{p_{2}})|<|N^{+}(u_{p_{2}})|,$
and $u_{p_{2}}$ is not a degree-doubling node.
Working backward, since this holds for all $u_{p_{2}}\in R_{i-1}$, the interior degree condition
$|\text{int}(u_{p_{2}},u_{i})|\geq i.$
must hold for all parents $u_{p_{2}}$ and children $u_{i}$. ∎
These two proofs keep the complexity of the algorithm linear. We do not need to check every node in the previous rooted neighborhood against every node in the current rooted neighborhood. Instead, we need to only have a representative from the previous rooted neighborhood. Lemma 8.2 (One Interior Succeed Means All Succeed) states that if the interior bound holds for that representative, then it will hold for the entire rooted neighborhood. What that means is that the node in $R_{i}$ is doing its part of the load balancing for all its parents. Lemma 8.1 (One Interior Fail Means All Fail) is the converse of this. It also takes a representative from the previous rooted neighborhood and checks the interior out-degree. If that number does not meet the requirement, then not only does the representative become a degree-doubling node, but every other member of that rooted neighborhood does. This happens because the second out-neighbors of that representative are also second members of the other members of that rooted neighborhood.
Algorithm 3 Decreasing Neighborhood Sequence Algorithm
1:Determine a minimum out-degree node $v_{0}\in G$.
2:Partition $V$ into ordered sets $R_{0},R_{1},\dots,R_{k}$ where
$R_{0}=\{v_{0}\},\quad R_{i}=N^{+}(R_{i-1})\setminus\bigcup_{j=0}^{i-1}R_{j}\quad\text{for }i>0.$
3:for $i=0$ to $k$ do
4:     for each node $u_{i}\in R_{i}$ do
5:         if there exists an arc $u_{i}\to w\in G$ with $w\in R_{m}$ where $m<i$ then
6:              Mark $u_{i}$ as degree_doub[back]
7:              HALT
8:         else
9:              if $i>0$ then
10:                  Let $u_{p}\in R_{i-1}$ be a representative parent of $u_{i}$.
11:                  if $|\text{int}(u_{p},u_{i})|<i$ then
12:                       Mark $u_{p}$ as degree_doub[dense]
13:                       HALT
14:                  end if
15:              end if
16:         end if
17:     end for
18:     if $i>0$ and $|R_{i}|\leq 2$ then
19:         Mark $u_{p}\in R_{i-1}$ as degree_doub[size]
20:         HALT
21:     end if
22:end for
[…]
Theorem 8.2.
(Algorithm Complexity) Suppose we have a graph $G$. The Decreasing Neighborhood Sequence Property Algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ represents the number of vertices in $G$ and $|E|$ represents the number of edges in $G$.
Proof.
We will first show that the complexity of building the partitions has a worst-case complexity of $O(|V|+|E|)$.
Given a graph $G$, the search for a minimum out-degree node that’s $O(|V|)$ nodes. Then, to evaluate the degrees of each of those nodes is a constant factor. This makes the search for the minimum out-degree node $O(|V|)$.
Once we have this value, we will place every node and every edge of $G$ into a pre-processing array ready to be partitioned. Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$. Each node will individually be processed based on its distance from the minimum out-degree node, along with the connected outgoing arcs. This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into neighborhoods. This means that the partitioning has a complexity of $O(|V|+|E|)$.
Next, we will look into the complexity of the run-time analysis of the Decreasing Sequence Algorithm.
First, remember that we proved in the Lemma 6.4 (Interior Degrees Doubles) that every neighborhood has a node whose interior degree doubles. This proof was actually the first instance where we saw that there were multiple degree doubling nodes. Every node in every cycle had its interior degree double. This lemma ensures that we don’t need to exhaustively check every node in $R_{i}$ for interior degree doubling; the existence of such a node is guaranteed, allowing us to potentially stop the algorithm once such a violation is found for a representative node, contributing to the linear complexity.
What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem. Solving the full problem requires determining the number of rooted neighborhoods necessary for this partition, which is bounded above by $O(|V|)$. The total number of rooted neighborhoods cannot exceed the number of nodes, $|V|$, since neighborhoods are based on partitions. Processing each rooted neighborhood’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$. This gives the run time a complexity of $O(|V|+|E|)$. ∎
This complexity analysis shows that the Decreasing Neighborhood Sequence Property Algorithm can effectively find a degree-doubling node in polynomial time. Moreover, this will work for any oriented graph. Complexity analysis shows that it is not hindered by larger graphs.
A further point is to consider the concept in Theorem 8.2 (Algorithm Complexity) of all nodes inside a neighborhood having their degrees double. It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
We opened by showing several applications of the SSNC, including social media, social network modeling, and network analysis. This highlights the necessity for an algorithm that can work efficiently in practice. This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph, is important. This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively. That makes it potentially applicable to real-world networks with millions or billions of nodes and edges. This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
Remark 8.1.
The structure imposed by the Decreasing Neighborhood Sequence Property does not just guarantee the existence of a degree doubling node. It does so with a progression of neighborhoods that decrease in size, culminating in dense neighborhoods that enforce multiple instances of degree-doubling. Thus, Seymour’s conjecture holds robustly under these conditions.
9 Applications
9.1 Network A/B Testing
The SSNC is not just a theoretical problem; it has practical applications. We have constructed an algorithm that searches for and finds nodes that satisfy this conjecture. Not only does our algorithm find these nodes, but it partitions the graph into a data structure, which can be beneficial for many other purposes. While the conjecture only asked for a single node, our approach finds a set of degree-doubling nodes, which are critical in understanding both network topology and influence dynamics.
By Theorem 8.2 (Degree-Doubling by Exterior Decreasing), we have developed an algorithm that identifies degree-doubling nodes. We then showed in Theorem 8.2 (Algorithm Complexity) that this algorithm has a complexity of $O(|V|+|E|)$. This makes it scalable to large graphs. In social networks with millions of users, the how fast the degree-doubling node identification algorithm is crucial. A less efficient algorithm would make it impractical to analyze the network and identify influential users within a reasonable time-frame.
A/B testing is often considered by companies when thinking of introducing a new product or service to customers. Users are randomly divided into two groups, a control group and a treatment group. It gives data driven decisions that focus on user feedback. Network A/B testing extends this methodology to social networking experiments. Users, here, are part of an interconnected system. Similar to traditional A/B tests, different versions of a post or an ad are shown to different audiences. Engagement metrics such as clicks or conversions are then compared. There are significant challenges to network A/B testing though. One such challenge is network interference, where users in the control group interact with users in the treatment group, unintentionally influencing the results. This limitation causes some researchers to dismiss network A/B testing entirely.[25]
Network A/B testing often relies on assumptions about types of influence, which can lead to misleading results. This includes homogeneous influence which assumes that all users have the same influence. A second type of influence is random influence, which says that influence is spread randomly throughout the network. [15] Finally, there is the assumption that users are independent of each other and that their actions are not influenced by the actions of others.
The Graph Level Order data structure is a way to organize an oriented graph into leveled neighborhoods. Nodes are placed based on their distance from a source node (normally a minimum out-degree node). Nodes are stored into ordered neighborhoods $R_{1}$, $R_{2}$, …, $R_{n}$, where $CR_{i}$ contains nodes that are distance $i$ from the root. Each neighborhood then connects to its next neighborhood.
The first thing this data structure does is helps with the problem of partitioning data sets. The data structure is defined with unique paths from $v_{0}$ to each hyper-node $R_{i}$. The hyper-nodes $R_{i}$ then connect only to the next hyper-nodes $R_{i+1}$. Thus, by the definition of the data structure, these sets $R_{i}$ and $R_{i+2}$ have no edges in common. This allows us to more easily set up the control group and the treatment group. The partitioning ensures that there can be no spill-over or network interference between the two groups. The Graph Level Order, by organizing the graph based on distance from a source node, helps to minimize network interference. Because nodes in different neighborhoods (e.g., $R_{i}$ and $R_{i+2}$) have no direct connections, the likelihood of users in the treatment group influencing users in the control group is significantly reduced. This allows for more accurate measurement of the treatment’s true effect.
The Graph Level Order, combined with our algorithm, allows for parallel processing of the graph. Because the neighborhoods in the Graph Level Order are independent, the computation of degree-doubling nodes within each neighborhood can be performed concurrently. The ability to parallelize the algorithm significantly reduces the computation time, especially for very large social networks. This makes it possible to analyze the network and identify influential users in real-time or near real-time, which is essential for dynamic network A/B testing.
Secondly, we have solved the SSNC with a set of degree-doubling nodes. These nodes are proven to be mathematically influential. These are not just intuition, but an answer to the question of ”which person has more followers post their social media information than they post themselves?” Thus, we can go back into the original social network, knowing these nodes are influential with more confidence about both our control and treatment sets. By identifying degree-doubling nodes, we can strategically select users for the A/B test. These nodes are likely to be highly influential within their communities. Including them in the treatment group allows us to observe how the new feature spreads through the network, providing a more realistic assessment of its potential impact. Alternatively, we might choose to exclude them from the treatment group to see how the feature performs without the amplification effect of these key influencers, giving us a different but equally valuable perspective.
Figure 13: Illustration showing partition of the Graph Level Order into two groups, ready for Control and Treatment.
This application of the SSNC seeks to plant the seeds for future research in network A/B testing. There are other issues that the Graph Level Order algorithm does not resolve. One limitation is that this algorithm will require an oriented graph, instead of a standard directed graph. that may limit applications as well. However, this research may hold a key to unlocking network A/B testing.
10 Conclusion
In this paper, we approached the SSNC from a different, contradictory, angle. This allowed us to come up with a data structure that well ordered the nodes of the graph into neighborhoods. The data structure gave rise to a linear time algorithm that solves the SSNC and finds degree-doubling nodes. This data structure also adds to the mathematical literature on social media by dissecting transitive triangles into six distinct types that can help the SSNC with friendship recommendations. Most of the paper has focused on the mathematical aspects of this problem, trying to rigorously and constructively justify these claims. There are many real-world applications of both the Graph Level Order and the algorithm. This was then shown to be applicable to network A/B testing, which makes the SSNC solution extremely relevant. Future work will look into investigating knowledge graphs for things like entities, emotions or hate speech in social media. This research is still ongoing and holds promise for understanding more about user behavior and particularly in the realm of natural language processing.
This exploration of the SSNC has allowed us to look through the lenses of graph theory, discrete mathematics, combinatorics, abstract algebra, algorithms, and data structures. This research helped us to use many mathematical tools, like both total orders, partitioning, divide and conquer, strong and weak induction, recurrence relations, and pathfinding techniques. One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the SSNC with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum out-degree node and can partition the graph into neighborhoods and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the SSNC. It was no longer framed as merely a graph theory problem. By utilizing the partition strategy and solving different parts independently, this algorithm shows that the SSNC could effectively be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The SSNC started as a question in 1990. Over the next 35 years, it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, it’s time to begin looking forward. The SSNC can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes will be infected people, and the arcs would be one-way infections. The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly? There are many more real-life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, and neural networks. There are too many to name in this short paper.
We have presented a linear time algorithm that can work on large graphs in these problems. By constructing the algorithm for the SSNC and proving that the algorithm will find a degree-doubling node, we have proved that the algorithm is theoretically sound.
Secondly, we introduced a data structure, the Graph Level Order (GLOVER). This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for the Graph Level Order—the well ordering and the dual metrics of distance and degrees. Still, the ability for the data structure to partition the nodes of an oriented graph into neighborhoods, which then allows for powerful methods like partitioning and mathematical induction, can have further applications. This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
Future research and development should explore other possibilities for this data structure as well as their complexity and scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
Thirdly, this research offered an alternative method for the SSNC, moving beyond existential proofs to a constructive, algorithmic solution. Instead of simply demonstrating the existence of a node whose out-degree doubles in the graph’s square, we focused on identifying such a node through a novel approach. This involved several key innovations.
The first of these techniques is graph partitioning. We dissected the problem by strategically partitioning the graph. This revealed underlying properties about these partitions and how they relate to one another that were crucial to our algorithmic solution. These properties of the partitions allowed us to analyze the relationships between a vertex and its second out-neighbors in a new light.
Moreover, we employed a dissection strategy after that partition. After breaking down the problem into smaller, more manageable subproblems, we solved them in an independent nature. This was akin to divide and conquer, but not quite since there was no need for a recurrence relation. Instead we simply solved each independent subproblem in $O(1)$ time. We were then able to return these independent subproblems back to the overall problem, which gave rise to a global solution. This approach, combined with our partitioning technique, enabled us to develop an efficient pathfinding algorithm.
Lastly, we introduced the concept of dual metrics. This allowed us to simultaneously consider multiple aspects of the problem simultaneously. Instead of making the metrics of degree and distance compete with one another, we utilized them both. Distance was selected as an outer metric to partition the nodes, while degree was selected as an inner metric to help differentiate the nodes within the rooted neighborhoods, and establish a total order. This multifaceted perspective proved essential in connecting the local degree properties of vertices within a neighborhood to the global neighborhood property of the distance required by the conjecture.
These techniques, particularly the graph partitioning and pathfinding strategies, may be applicable to other problems in graph theory and related domains. While the specific degree and distance dual metrics used in the SSNC proof might not be directly transferable, the general principle of considering multiple perspectives could be valuable. It is possible that other problems will require the development of new data structures and metrics tailored to their specific characteristics. However, the core idea of strategically dissecting the problem, as demonstrated in our solution to the SSNC, offers a promising direction for future research.
This shift in emphasis from existential proofs to algorithmic solutions, combined with the development of new data structures and analytical techniques, may encourage researchers to explore constructive approaches with a focus on practical applications. We hope that this work will inspire further research on such approaches to other graph theory problems. We also encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
11 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:5-6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, wife and children—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
I also appreciate the Department of Mathematics at Morehouse College for helpful discussions that were valuable during the research process. In particular, I thank Dr. Duane Cooper for asking the clarifying question where he asked for an example showing that Conjecture 1.1 and Conjecture 3.1 were equivalent.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The Graph Level Order is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
Generative AI usage in the review process: Both ChatGPT 3.5 (July 20, 2024 version, published by OpenAI, https://chat.openai.com/) and Gemini (2.5 Flash version, Published by Google, https://gemini.google.com/) were used to analyze text drafts of human writing. All AI output was evaluated and judged based on merit by a human before being included anywhere in the manuscript.
Appendix A Appendix 1: Definitions
Definition A.1.
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(u,u)$ where $u$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $(u,v)$ and $(v,u)$ where $u$ and $v$ are nodes in $G$.
Definition A.2.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of arcs defined as:
$E^{2}=\{(u,v)\mid(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E\}$
Definition A.3.
The distance between nodes $u$ and $v$, denoted $\text{dist}(u,v)$, is the length of the shortest directed path from $u$ to $v$.
Definition A.4.
Let $G=(V,E)$. The first out-neighborhood of a vertex $v\in V$ is defined as:
$N^{+}(v)=\{w\in V\mid(v,w)\in E\}$
Definition A.5.
Let $G=(V,E)$. The second out-neighborhood of a vertex $v\in V$ is defined as:
$N^{++}(v)=\{u\in V\mid\exists w\in V\text{ such that }(v,w)\in E\text{ and }(w,u)\in E,\text{ and }u\notin N^{+}(v)\}$
Definition A.6.
Let $G=(V,E)$. Let $S\subseteq V$ be a subset of the vertices of G. Then the induced subgraph $G[S]$ is the graph whose vertex set is $S$ and whose edge set consists of all the edges in $E$ that have both endpoints in $S$.
Definition A.7.
In an oriented graph $G=(V,E)$, a node $v\in V$ is a degree-doubling node (or Seymour vertex) if $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}(v)$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$ in $G$, respectively.
Definition A.8.
A rooted neighborhood $R_{i}$ is the subgraph of $G$ induced by the set of nodes at distance $i$ from $v_{0}$, given a minimum out-degree node $v_{0}$. Formally,
$R_{i}=(V_{i},A_{i})$
where
$V_{i}=\{u\in V(G):\text{dist}(v_{0},u)=i\}$
and
$A_{i}=\{(u,v)\in E(G):u,v\in V_{i}\}$
where $V(G)$ and $E(G)$ denote the graph $G$’s vertex and edge sets, respectively, and $\text{dist}(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Definition A.9.
Let $(x,y),(x,u),(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle, where $x$ is a common predecessor of both $y$ and $u$, and $y$ also connects to $u$.
Definition A.10.
Let $u_{i}$ be a node in the rooted neighborhood $R_{i}$ for some $i\geq 0$. A child of $u_{i}$ is a node $v_{i+1}\in R_{i+1}$ such that $(u_{i},v_{i+1})\in G$ (we also say that $u_{i}$ is the parent of $v_{i+1}$).
Definition A.11.
Interior Neighbor and Interior Degree Let $u_{i}\in R_{i}$ be a parent node with children $v_{1},v_{2}\in R_{i+1}$. We define the interior neighbors of $v_{1}$ with respect to $u_{i}$ as those nodes $z\in V$ such that both $(u_{i},z)\in E$ and $(v_{1},z)\in E$. That is, nodes that are common out-neighbors of both $u_{i}$ and $v_{1}$, forming transitive triangles.
$\text{int}(u_{i},v_{1}):=N^{+}(u_{i})\cap N^{+}(v_{1})$
The interior degree of $v_{1}$ with respect to $u_{i}$, denoted $deg_{int}(u_{i},v_{1})$ is defined as:
$deg_{\text{int}}(u_{i},v_{1}):=|\text{int}(u_{i},v_{1})|$
Definition A.12.
Let $u_{i}\in R_{i}$ be the parent of $v,w\in R_{i+1}$. Then $v$ and $w$ are said to be siblings.
Definition A.13.
Let $u_{i}\in R_{i}$ be a parent of a node $v_{i+1}\in R_{i+1}$. The exterior neighbors of $v_{i+1}$ with respect to $u_{i}$ are nodes $z$ such that $z$ is a second out-neighbor of $u_{i}$ and a first out-neighbor of $v_{i+1}$, i.e., $z\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$. This implies that there exists a path $u_{i}\rightarrow w\rightarrow z$, and an arc $v_{i+1}\rightarrow z$ exists, but there is no direct arc $u_{i}\rightarrow z$. Unlike the interior neighbors, exterior neighbors are neighbors of the child that are not shared by the parent.
The exterior degree of $v_{i+1}$ with respect to $u_{i}$ is $|\text{ext}(u_{i},v_{i+1})|$.
Definition A.14.
Let $v_{0}$ be a minimum out-degree node. Suppose that $x$ is a node in the rooted neighborhood $R_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in R_{j}$, where $j<i$.
Definition A.15.
A Graph Level Order (GLOVER) on a directed graph $G=(V,E)$ can be defined as follows:
  1. 1.
    Leveled Rooted Neighborhood Structure: The vertices of $V$ with minimum out-degree node $v_{0}$ are partitioned into levels of Rooted Neighborhoods $R_{1},R_{2},\ldots,R_{n}$, where $R_{i}=\{v\in V:\text{dist}(v_{0},v)=i\}$, and $\text{dist}(v_{0},v)$ is the shortest path from $v_{0}$ to $v$.
  2. 2.
    Universal Rooted Neighborhood Order: The rooted neighborhoods are totally ordered such that $R_{i}<R_{j}$ if and only if $i<j$.
  3. 3.
    Comparability Within Rooted Neighborhoods: For any two vertices $u,v\in R_{i}$, their order is determined based on a specific metric (e.g., degree).
  4. 4.
    Universal Vertex Order: For any two vertices $u\in R_{i}$ and $v\in R_{j}$ with $i<j$, $u$ is considered less than $v$.
  5. 5.
    Interior and Exterior Out-neighbors: For a node $u\in R_{i}$ and $v\in R_{i+1}$, where $u$ is the parent of $v$
    • The interior neighbors of $u$ and $v$ are defined by the set $\text{int}(u,v)$.
    • The exterior neighbors of $u$ and $v$ are defined by the set $\text{ext}(u,v)$.
Definition A.16.
For a node $u\in G$ in an oriented graph $G$, we say that $u$ has the Decreasing Neighborhood Sequence Property if the size of its first out-neighbors is strictly larger than the size of its second out-neighbors, i.e., $|N^{++}(u)|<|N^{+}(u)|$.
Definition A.17.
Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
Definition A.18.
Let $x\in R_{i}$. Define $\text{int}(R_{i-1},x)$ as the first out-neighbors of $x$ within $R_{i}$, and $\text{int}^{++}(R_{i-1},x)$ as the second out-neighbors of $x$ within $R_{i}$. We say that $x$ has its interior degree doubled if $|\text{int}^{++}(R_{i-1},x)|\geq|\text{int}(R_{i-1},x)|$.
Appendix B Appendix 2: Graph Level Order Applications
The deciding factor that held the algorithm together and helped solve the SSNC was the Graph Level Order data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs—partitioning, ordering those partitions, and adding interior and exterior arcs—until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the SSNC, exterior arcs do not always have to only be defined to the next neighborhood. That was a problem-specific definition.
Moving forward, the Graph Level Order should be well suited to tackle problems in graph theory and beyond. By introducing neighborhoods, another way of doing induction on graphs has been introduced. The concept of partitioning, which has dominated much of computer science, shows up in this paper as well. We can treat these neighborhoods independently of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the Graph Level Order is moving forward.
What makes Graph Level Order special are things like a total order on neighborhoods. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
B.1 Representation
The more popular representations of graphs are adjacency matrices, array-lists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a Graph Level Order, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be a new level of data: rooted neighborhoods, to go along with the standard data serialization. The neighborhoods will be given an ID associated with their distance from the minimum out-degree node. The nodes will then be assigned to their unique neighborhoods. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
Node Targets Neighborhood
0 1, 2, 3 $R_{0}$
1 2, 4, 5, 6 $R_{1}$
2 3, 4, 5, 6 $R_{1}$
3 1, 4, 5, 6 $R_{1}$
4 5, 7, 8 $R_{2}$
5 6, 7, 8 $R_{2}$
6 7, 8, 1 $R_{3}$
7 8, 6, 1 $R_{3}$
8 6, 7, 1 $R_{3}$
Table 5: This figure illustrates a JSON representation of a Graph Level Order in array-list representations. Each node has a list of targets, along with its assigned neighborhood. Not shown are the nodes in neighborhood $R_{4}$. In this JSON format, each key represents a node ID, and its value is an object containing ’targets’ (a list of its out-neighbors) and ’neighborhood’ (the rooted neighborhood it belongs to, $R_{0}$ being the minimum out-degree node). This structure explicitly groups nodes by their distance from $v_{0}$.
The determination of the number of rooted neighborhoods and the assignment of nodes to rooted neighborhoods also do not need to be declared beforehand. Instead, just as the data structure uses the minimum function to find a node representing the minimum out-degree node $v_{0}$. This same function that determines the minimum out-degree node will return a value $\delta=d^{+}(v_{0})$, and $\delta$ will represent the maximum number of rooted neighborhoods. We would have a similar function to assign nodes to rooted neighborhoods. This can also give us a serialization.
With this rooted neighborhood representation of the SSNC, by way of the Graph Level Order, we can begin to visualize elements of these lemmas. For example, Lemma 6.8 (DNSP Neighborhood Size Property) can be visualized by Example 10. This shows a series of ordered rooted neighborhoods. Each rooted neighborhood has fewer nodes. The fewer nodes imply that the rooted neighborhood circle can be drawn with a smaller radius, giving light to the decrease in size.
Next, these nodes are not just treated like clusters inside these neighborhoods but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth-first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
Figure 14: Here we have an illustration of some of the most common proofs in mathematics. They are partitioned by proof types. Interior arcs are drawn between nodes in the same neighborhood, representing things proven the same way. Exterior arcs are drawn between nodes that can be proven through multiple methods.
The Graph Level Order extends beyond oriented graphs though. This data structure can be used to represent arbitrary JSON/YAML/XML datasets. Anything that can be represented in one of these can be represented by a Graph Level Order and will have the tools of Graph Theory at hand.
Theorem B.1.
Any dataset that can be represented in a structured encoding format (such as JSON, XML, YAML, or equivalent) can be represented using a Graph Level Order, provided the user selects a valid metric. This will yield a valid Graph Level Order structure.
Proof.
Structured encoding formats, including JSON, XML, and YAML, encode data as discrete objects with finite fields and elements. Consequently, the number of elements in any dataset represented by these formats is finite, ensuring the data can be fully enumerated and partitioned.
Let $M$ be the metric chosen by the user such as degree or distance in the SSNC. $M$ is a field, attribute, or key present in the encoded data. $M$ corresponds to the basis for defining relationships across the neighborhoods.
For each element $x$ in the dataset, the metric $M(x)$ determines a value or set of values that can be used to group elements. The dataset is partitioned into neighborhoods $R_{1},R_{2},...,R_{k}$ based on $M$, where $R_{i}=\{x\mid M(x)=v_{i}\}$ for some value $v_{i}$.
If $M$ improperly defines neighborhoods (e.g., allows elements to belong to multiple neighborhoods or leaves elements unassigned), then $M$ is not a valid metric. A valid $M$ ensures that all elements are correctly assigned to exactly one neighborhood.
Thus, any dataset encoded in a structured format can be represented by a Graph Level Order, provided a suitable metric $M$ exists. ∎
The thing about the Graph Level Order is the adaptability. It adapts to the environment, and in particular the partition of that environment. SSNC called for a partition that needed an anchor node. In general most partitions do not require this and so we will not be seeing these types of graphs in the future unless especially called for. What we will be seeing are requests for partitions. Numeric partitions like distance, generally give rise to orderings easier. Other partitions may give rise to lexicographic orderings. However, the largest benefit of the Graph Level Order moving forward is extending graph theory concepts beyond standard graph theory and into the world of JSON/YAML/XML.
Appendix C Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second out-neighbors or squares of graphs are probably not either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the SSNC led this paper into the terms Decreasing Neighborhood Sequence Property, induced subgraphs, neighborhoods, interior out-neighbors, exterior out-neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book ”Philosophy of Mathematics” speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand the research. It also has things like MathJax installed to read lemmas, definitions, and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
Readers are encouraged to explore the website for additional information and examples at
https://glovermethod.com.
This is an evolving project. Additional examples and applications will continue to be added.
Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.